### v2

# Bivariate Matrix-valued Linear Regression (BMLR): Finite-sample performance under Identifiability and Sparsity Assumptions

Abstract
This study explores the estimation of parameters in a matrix-valued linear regression model, where the $T$ responses $(Y_{t})_{t=1}^{T}\in\mathbb{R}^{n\times p}$ and predictors $(X_{t})_{t=1}^{T}\in\mathbb{R}^{m\times q}$ satisfy the relationship $Y_{t}=A^{*}X_{t}B^{*}+E_{t}$ for all $t=1,\ldots,T$. In this model, $A^{*}\in\mathbb{R}_{+}^{n\times m}$ has $L_{1}$-normalized rows, $B^{*}\in\mathbb{R}^{q\times p}$, and $(E_{t})_{t=1}^{T}$ are independent noise matrices following a matrix Gaussian distribution. The primary objective is to estimate the unknown parameters $A^{*}$ and $B^{*}$ efficiently.
We propose explicit optimization-free estimators and establish non-asymptotic convergence rates to quantify their performance. Additionally, we extend our analysis to scenarios where $A^{*}$ and $B^{*}$ exhibit sparse structures. To support our theoretical findings, we conduct numerical simulations that confirm the behavior of the estimators, particularly with respect to the impact of the dimensions $n,m,p,q$, and the sample size $T$ on finite-sample performances. We complete the simulations by investigating the denoising performances of our estimators on noisy real-world images.
1 Introduction
Supervised learning is frequently applied to large databases and with the recent advancements in data collection techniques, matrix structures have become increasingly prevalent in collected data. Examples include images, longitudinal multivariate measurements, and spatiotemporal data, where rows and columns often represent distinct dimensions, such as time, space, or experimental conditions. Consequently, regression tasks now often involve matrix-structured responses using matrix-structured predictors. This shift introduces both opportunities and challenges, as traditional methods designed for scalar or vector responses are often inadequate for preserving the inherent structure of matrix-valued data.
Matrix-valued models for regression tasks have been extensively studied in a setting where rows correspond to vector-valued observations (Bunea et al.,, 2011; Giraud,, 2011; Bettache & Butucea,, 2023), particularly as random matrix theory has developed strong theoretical foundations (Vershynin,, 2010; Tao,, 2012). Matrix-valued models are also extensively studied in topic models and other matrix factorization procedures (Bing et al., 2020c, ; Bing et al., 2020b, ; Bing et al., 2020a, ; Eisenach et al.,, 2020; Jin et al.,, 2021; Bing et al., 2022b, ; Bing et al., 2022a, ; Bunea et al.,, 2022; Ke & Wang,, 2024). More recently, there has been growing interest in matrix-valued models designed specifically for matrix-valued observations. These models provide a powerful framework for capturing complex relationships between predictors and responses in regression settings (Li et al.,, 2010; Hoff,, 2015; Ding & Cook,, 2018; Kong et al.,, 2020; Chen & Fan,, 2023; Chen & Tang,, 2024; Boyle et al.,, 2024). Similarly, in autoregression, they address dynamic dependencies within matrix-structured time series (Chen et al.,, 2021, 2022; Yu et al., 2024a, ; Yu et al., 2024b, ; Jiang et al.,, 2024; Tsay,, 2024; Bettache,, 2024; Lam & Cen,, 2024; Samadi & Billard,, 2024; Zhang et al.,, 2024). By leveraging the structure of matrix-valued data, these approaches not only enhance interpretability but also improve predictive performance.
This work investigates the Bivariate Matrix-Valued Linear Regression (BMLR) model under identifiability conditions. We introduce optimization-free estimators and derive non-asymptotic upper bounds on their convergence rates. Notably, this is the first study to provide a comprehensive non-asymptotic analysis for each parameter individually, filling a critical gap in the existing literature.
1.1 Model
Consider $T$ observations $(X_{t},Y_{t})_{t=1}^{T}$ satisfying the model:
$Y_{t}=A^{*}X_{t}B^{*}+E_{t},\quad t\in[T]:=\{1,\ldots,T\}$ (1)
where matrices $(X_{t})_{t=1}^{T}\in\mathbb{R}^{m\times q}$ are the predictors, the parameter $A^{*}\in\mathbb{R}_{+}^{n\times m}$ is an unknown matrix that, for identifiability reasons, is assumed to have rows with unit $L_{1}$ norm and the parameter $B^{*}\in\mathbb{R}^{q\times p}$ is another unknown matrix. For statistical inference we assume that the noise matrices $(E_{t})_{t=1}^{T}\in\mathbb{R}^{n\times p}$ are independent noise matrices following a matrix normal distribution (Dawid,, 1981; Waal,, 2004; Gupta & Nagar,, 2018; Barratt,, 2018; Mathai et al.,, 2022):
[…]
where $\Sigma_{r}\in\mathbb{R}^{n\times n}$ is the row-covariance matrix and $\Sigma_{c}\in\mathbb{R}^{p\times p}$ is the column-covariance matrix, both assumed to be known. The objective is to estimate the matrices $A^{*}$ and $B^{*}$ based on the observed $(X_{t},Y_{t})_{t=1}^{T}$, allowing to predict the unobserved value $Y_{t+1}$ given an out-of-sample $X_{t+1}$.
We call (1) the bivariate matrix-valued linear regression (BMLR) model. Notice that for any $t\in\left[T\right]$, $A^{*}$ captures the link between the rows of $Y_{t}$ and the rows of $X_{t}B^{*}$ while $B^{*}$ captures the link between the columns of $Y_{t}$ and the columns of $A^{*}X_{t}$. This model is closely related to the standard multivariate linear regression model. If the matrix structures of the responses and predictors are ignored, the model is equivalent to
$\text{vec}(Y_{t})=((B^{*})^{\top}\otimes A^{*})\text{vec}(X_{t})+\text{vec}(E_{t}),$
where $\text{vec}(Y_{t})$ denotes the vector obtained by stacking the columns of $Y_{t}$ (Henderson & Searle,, 1981). This procedure can be treated as a standard multivariate linear regression model with $T$ observations. One would be able to compute $\hat{M}$, an estimator of $M^{*}=(B^{*})^{\top}\otimes A^{*}$ (Chen & Tang,, 2024). However, it would not be possible to estimate $A^{*}$ and $B^{*}$ together with their matrix structure from this procedure. Indeed the nearest Kronecker product, aiming to estimate $\hat{A}$ and $\hat{B}$ from $\hat{M}$, is equivalent to finding a nearest rank-1 matrix (Loan,, 2000). Hence the matrix attributes of $A^{*}$ and $B^{*}$ are ignored in this procedure.
1.2 Context
Given the observations $(X_{t},Y_{t})_{t=1}^{T}$, the linear regression framework models the relationship between the responses and the predictors using an unknown linear map $f:\mathbb{R}^{m\times q}\rightarrow\mathbb{R}^{n\times p}$. For each $t\in\left[T\right]$ this relationship is expressed as $Y_{t}=f(X_{t})+E_{t}$ where $E_{t}$ represents the noise term. The goal is to learn the unknown function $f$ only using the given observations. In a parametric setting we assume that $f$ belongs to a predefined class of functions, typically parameterized by a finite number of parameters.
A naive approach would be to break down the linear relationship by focusing on each entry of the responses individually. This leads to consider for $i\in[n]$ and $j\in[p]$ a linear functional $f_{ij}:\mathbb{R}^{m\times q}\rightarrow\mathbb{R}$ such that
$[Y_{t}]_{ij}=f_{ij}(X_{t})+[E_{t}]_{ij},\quad t\in\left[T\right],$
where $[Y_{t}]_{ij}$ denotes the coefficient on the $i^{th}$ row and $j^{th}$ column of the matrix $Y_{t}$. Here, Riesz representation theorem (Brezis,, 2011) ensures the existence, for fixed $i\in[n]$ and $j\in[p]$, of a unique matrix $M^{*}_{ij}\in\mathbb{R}^{m\times q}$ such that
$[Y_{t}]_{ij}=\text{Tr}\left(X_{t}^{\top}M^{*}_{ij}\right)+[E_{t}]_{ij},\quad t\in\left[T\right],$
where $X_{t}^{\top}$ stands for the transposed of $X_{t}$ and $\text{Tr}\left(X_{t}^{\top}M^{*}_{ij}\right)$ denotes the trace of the square matrix $X_{t}^{\top}M^{*}_{ij}$, which is the canonical scalar product $\langle X_{t},M^{*}_{ij}\rangle_{F}$. In this model, the objective is to estimate the $np$ matrices $(M^{*}_{ij})$. This is a parametric problem which is formally of dimension $npmq$. If the matrices $M^{*}_{ij}$ are assumed to be of rank $r\ll\min(m,q)$, the intrinsic dimension of the problem is reduced to $npr(m+q-r)\approx npr\max(m,q)$. This problem is equivalent to considering $np$ independent trace regression models (Rohde & Tsybakov,, 2011). Hence, this naive approach ignores the multivariate nature of the possibly correlated entries of each response $Y_{t}$.
To overcome this issue, we consider for $i\in[n]$ and $j\in[p]$ the vectors $\alpha_{i}\in\mathbb{R}^{m}_{+}$ and $\beta_{j}\in\mathbb{R}^{q}$ and assume $M^{*}_{ij}=\alpha_{i}\beta_{j}^{\top}$. For identifiability issues we assume that for all $i\in[n]$, $\|\alpha_{i}\|_{1}=1$. This model ensures that the matrices $(M^{*}_{ij})$ share a common structure. It also implies that they share the same rank, namely one here. Hence this model now accounts for the multivariate nature of the problem. It has $n(m-1)+pq$ free parameters and rewrites as follows:
$[Y_{t}]_{ij}=\text{Tr}\left(X_{t}^{\top}\alpha_{i}\beta_{j}^{\top}\right)+[E_{t}]_{ij},\quad t\in\left[T\right].$
Consider $A^{*}$ the matrix obtained by stacking the $n$ row vectors $\alpha_{i}$ and $B^{*}$ the matrix obtained by stacking the $p$ column vectors $\beta_{j}$. This leads to the BMLR model (1).
1.3 Related works
The BMLR model, first introduced in (Hoff,, 2015), has gained notable attention as a powerful framework for examining relationships between matrix-structured responses and predictors. In the development of estimation techniques for this model, two principal methods have been explored: alternating minimization and spectral aggregation.
Alternating Minimization, usually presented in a Maximum Likelihood Estimation (Ding & Cook,, 2018) or a least squares Estimation(LSE) context (Hoff,, 2015). As the objective is non convex in both parameters, a two-step iterative algorithm is usually derived to construct the estimators.
Spectral Methods, presented in a factor model framework in (Chen & Fan,, 2023), offers an alternative that leverages the spectral properties of the target matrices. Authors propose a new estimation method, the $\alpha$-PCA, that aggregates the information in both first and second moments and extract it via a spectral method. They show that for specific values of the hyperparameter $\alpha$, namely $\alpha=-1$, the $\alpha$-PCA method corresponds to the least squares estimator. However, the procedure is non convex and they rely on approximate solution by alternating minimization. More specifically they maximize row and column variances respectively after projection.
Kronecker Product Factorization: The Kronecker Product Factorization (KRO-PRO-FAC) method has recently introduced new possibilities for estimation within matrix-valued linear regression. Chen & Tang, (2024) present this approach, which leverages Kronecker products to decompose complex matrices into simpler components. A key advantage of this method is that it circumvents the need to estimate the covariance between individual entries of the response matrix, significantly reducing computational complexity. The KRO-PRO-FAC algorithm is accompanied by non-asymptotic bounds on the estimation of the Kronecker product between the parameters. However, an important limitation of this approach is the lack of direct control over the estimation accuracy of each parameter separately. This restricts its applicability in scenarios where parameter-wise interpretability or precision is critical, leaving a gap that motivates alternative methodologies, such as the optimization-free estimators proposed in this work.
Autoregressive Frameworks: In autoregressive settings, where $X_{t}:=Y_{t-1}$ in (1), the primary focus is on capturing temporal dependencies by minimizing the residual sum of squares (Chen et al.,, 2021; Yu et al., 2024a, ; Yu et al., 2024b, ; Jiang et al.,, 2024; Tsay,, 2024). These frameworks are particularly relevant for modeling dynamic systems and matrix-valued time series. However, parameter estimation in autoregressive frameworks typically relies on computationally intensive procedures, such as iterative optimization or matrix decompositions. These methods become increasingly impractical as the dimensions of the data grow, creating a significant bottleneck in high-dimensional settings. This challenge is especially pronounced in applied studies, where the scale and complexity of datasets continue to expand, underscoring the need for more efficient and scalable estimation techniques.
1.4 Summary of Contributions
This work studies the estimation problem of the matrix parameters $A^{*}$ and $B^{*}$ in (1). Our contributions are the following:
  • Noiseless Case Analysis: In Section 2, we study the noiseless case and establish that, in the absence of noise, the true parameters can be exactly retrieved. This analysis highlights the fundamental identifiability properties of the model and provides a baseline for estimating the parameters in the noisy regime.
  • Optimization-Free Estimators: In Section 3, we propose explicit, optimization-free estimators $\hat{B}$ and $\hat{A}$ defined in (8) and (10) respectively. They significantly simplify the estimation process and are particularly advantageous in high-dimensional settings where traditional optimization-based methods become computationally prohibitive. Theoretical guarantees are provided in Theorems 3.4 and 3.7. We establish non-asymptotic bounds characterizing the dependence of estimation accuracy on the problem dimensions ($n,p,m,q$) and the sample size $T$. For $\hat{A}\in\mathbb{R}^{n\times m}$, the performance improves with larger sample size $T$ and larger values of $p$ and $q$, the column dimensions of the response and predictor matrices, respectively. However, the performance deteriorates as $n$ and $m$, which determine the size of $A^{*}$, increase. In contrast, for $\hat{B}\in\mathbb{R}^{q\times p}$, the convergence rate improves with increases in $T$ and $n$, showcasing a ”blessing of dimensionality” effect in the row dimension of the target matrices. Nonetheless, the performance decreases with larger values of $p$, $q$, and $m$.
  • Sparse Adaptive Estimators: Extending the framework to sparse settings, we introduce hard-thresholded estimators, $\hat{B}^{S}$ and $\hat{A}^{S}$, defined in (11) and (12) respectively. These estimators exploit the sparsity of the true parameters to achieve improved convergence rates, as established in Theorems 3.8 and 3.9. Moreover, we demonstrate that these estimators can recover the exact support of the true parameters with high probability, providing strong practical guarantees. These estimators exhibit the same dependency on the problem dimensions ($n,p,m,q$) and the sample size $T$ as their dense counterpart $\hat{B}$ and $\hat{A}$.
  • Numerical Validation: In Section 4, we validate our theoretical findings through extensive numerical simulations. On synthetic datasets, we show that the empirical convergence rates align closely with theoretical predictions, as illustrated in Figures 1 and 2. For real-world data from the CIFAR-10 dataset, we demonstrate the practical effectiveness of our framework by introducing controlled noise, estimating the correction matrices $\hat{A}$ and $\hat{B}$ on the training set, and evaluating denoising performance on the test set. The results, presented in Figures 3 and 4, highlight the effectiveness of the proposed estimators in practical scenarios.
Overall, this work offers a comprehensive framework that integrates theoretical rigor, computational efficiency, and practical applicability for matrix parameter estimation in high-dimensional settings.
2 Analysis in the Noiseless Setting
In this section we focus on the noiseless case of (1). We consider $T$ observations $(X_{t},M_{t})$ that satisfy the relationship
$M_{t}=A^{*}X_{t}B^{*},\quad t\in\left[T\right],$ (2)
where $A^{*}\in\mathbb{R}_{+}^{n\times m}$ and $B^{*}\in\mathbb{R}^{q\times p}$ are the unknown parameters to be recovered. The matrix $A^{*}$ is assumed to have rows with unit $L_{1}$-norm, ensuring identifiability of the model. We refer to the setup in (2) as the noiseless bivariate matrix-valued linear regression (NBMLR) model.
NBMLR provides a clean framework to study the fundamental properties of matrix recovery without the complications introduced by noise. Understanding the noiseless case is crucial, as it establishes baseline results that guide the analysis of the more realistic setting (1) involving noise. The primary goal here is to derive conditions under which the matrices $A^{*}$ and $B^{*}$ can be uniquely recovered, as well as to propose efficient algorithms for their recovery. This involves leveraging the structure of the observed predictors $(X_{t})_{t=1}^{T}$ and the constraints on the parameters.
2.1 Design Matrices form the Canonical Basis of $\mathbb{R}^{m\times q}$
To gain a deeper understanding of the underlying phenomenon, we first analyze a simplified case where the design matrices $(X_{t})_{t=1}^{T}$ are the elements of the canonical basis of $\mathbb{R}^{m\times q}$. In this setup we are given $T=mq$ observations $(X_{t},M_{t})_{t=1}^{T}$ where each predictor $X_{t}$ is a matrix with all entries set to zero except for a single entry equal to one. This setting is well studied in the vector case, typically referred to as a Sequence Model. Such models have been widely explored in the literature (Tsybakov,, 2009; Johnstone,, 2011). By focusing on this simple scenario, we aim to uncover insights that can extend to more complex cases.
Formally, for $(k,l)\in[m]\times[q]$, let $M_{(k,l)}$ corresponds to the basis element $E_{(k,l)}$, where $E_{(k,l)}$ is the matrix with all entries equal to zero except for a one in the $k$-th row and $l$-th column. The NBMLR model can be expressed as:
$M_{(k,l)}=A^{*}E_{(k,l)}B^{*},$ (3)
which corresponds to (2) for $t=k+(l-1)m$. The matrix $E_{(k,l)}$ acts as a selector, isolating the effect of the $k$-th row of $A^{*}$ and the $l$-th row of $B^{*}$.
Lemma 2.1.
In the NBMLR model (3), the parameters $A^{*}$ and $B^{*}$ can be explicitly recovered from the observed matrices $M_{(k,l)}$ as follows:
$\forall(l,j)\in[q]\times[p]:\quad[B^{*}]_{lj}$ $=\sum\limits_{k=1}^{m}[M_{(k,l)}]_{ij},\quad\text{for any }i\in[n],$
$\forall(i,k)\in[n]\times[m]:[A^{*}]_{ik}$ $=\dfrac{[M_{(k,l)}]_{ij}}{[B^{*}]_{lj}}\quad\text{for any }(l,j)\in[q]\times[p],\;\text{such that }[B^{*}]_{lj}\neq 0.$
2.2 Design Matrices Span $\mathbb{R}^{m\times q}$
In this section, we assume that the design matrices $(X_{t})_{t=1}^{T}$ form a generating family of $\mathbb{R}^{m\times q}$, which implies that $T$ is larger than $mq$. Notably, we no longer assume that the matrices $(X_{t})_{t=1}^{T}$ are elements of the canonical basis of $\mathbb{R}^{m\times q}$, nor do we require orthogonality among them.
We observe the matrices $(X_{t},M_{t})_{t\in[T]}$ following the noiseless model (2). To facilitate the analysis, we define two matrices, $\mathbb{M}\in\mathbb{R}^{T\times np}$ and $\mathbb{X}\in\mathbb{R}^{T\times mq}$ as follows
$\mathbb{M}:=\left(\text{vec}(M_{1}),\ldots,\text{vec}(M_{T})\right)\quad\text{and}\quad\mathbb{X}:=\left(\text{vec}(X_{1}),\ldots,\text{vec}(X_{T})\right),$ (4)
where $\text{vec}(\cdot)$ denotes the vectorization operator that transforms a matrix into a vector by stacking its columns.
Remark 2.2.
When the design matrices $(X_{t})_{t=1}^{T}$ are generated randomly, $\mathbb{X}^{\top}\mathbb{X}$ is invertible under mild conditions (Davidson & Szarek,, 2001; Mehta,, 2004; Rudelson,, 2008; Rudelson & Vershynin,, 2014).
We define the unobserved matrix $\mathbb{C}\in\mathbb{R}^{mq\times np}$ as
$\mathbb{C}:=\left(\text{vec}(A^{*}E_{(k,l)}B^{*})\right)_{k\in[m],\;l\in[q]},$
where each row of $\mathbb{C}$ corresponds to the vectorized form of $A^{*}E_{(k,l)}B^{*}$. The entry of $\mathbb{C}$ located at the $k+(l-1)m$-th row and $i+(j-1)n$-th column is denoted by $\left[\mathbb{C}\right]_{(k,l),(i,j)}.$
By construction, each entry of the matrix $\mathbb{C}$ is defined as:
$\left[\mathbb{C}\right]_{(k,l),(i,j)}=[A^{*}]_{ik}\cdot[B^{*}]_{lj},$
for all $i\in[n]$, $k\in[m]$, $l\in[q]$, and $j\in[p]$. This relationship ensures that:
$\mathbb{C}=B^{*}\otimes A^{*\top},$
where $\otimes$ denotes the Kronecker product.
The following lemma addresses an important question: in the NBMLR model defined by (2), can the matrix $\mathbb{C}$ be exactly reconstructed from the observations?
Lemma 2.3.
In the NBMLR model (2), where the design matrices $X_{t}$ form a generating family of $\mathbb{R}^{m\times q}$, the unobserved matrix $\mathbb{C}\in\mathbb{R}^{mq\times np}$ satisfies the following equality:
$\mathbb{C}=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}.$
Here, $\mathbb{M}\in\mathbb{R}^{T\times np}$ and $\mathbb{X}\in\mathbb{R}^{T\times mq}$ are defined in (4).
This result demonstrates that the matrix $\mathbb{C}$ can be computed directly from the observations. Moreover, it provides the foundation for recovering the unknown parameter matrices $A^{*}$ and $B^{*}$ exactly, as detailed in the subsequent discussion.
Proposition 2.4.
In the NBMLR model (2), where the design matrices $(X_{t})_{t=1}^{T}$ form a generating family of $\mathbb{R}^{m\times q}$, the parameter matrices $A^{*}\in\mathbb{R}^{n\times m}$ and $B^{*}\in\mathbb{R}^{q\times p}$ satisfy the following relationships:
$\forall(l,j)$ $\in[q]\times[p]:\quad[B^{*}]_{lj}=\sum_{k=1}^{m}\left[\mathbb{C}\right]_{(k,l),(i,j)},\quad\text{for any }i\in[n],$
$\forall(i,k)$ $\in[n]\times[m]:\quad[A^{*}]_{ik}=\frac{\left[\mathbb{C}\right]_{(k,l),(i,j)}}{[B^{*}]_{lj}},\quad\text{for any }(l,j)\in[q]\times[p]\;\text{such that }[B^{*}]_{lj}\neq 0.$
Remark 2.5.
We note that for fixed $l\in[q]$ and $j\in[p]$, the entries $\left(\left[\mathbb{C}\right]_{(k,l),(i,j)}\right)_{(i,k)}$ share the same sign. This arises from the bilinear structure of $\mathbb{C}$, as each entry is the product of $[A^{*}]_{ik}$ which is non-negative and $[B^{*}]_{lj}$
The following corollary provides a representation of the entries of $A^{*}$ and $B^{*}$ as averages. This characterization is particularly useful in noisy settings, as it facilitates the construction of estimators with significant variance reduction.
Corollary 2.6.
Let $\mathcal{D}\subset[p]\times[q]$ denote the set of indices $(j,l)\in[p]\times[q]$ such that $[B^{*}]_{lj}\neq 0$. Then the entries of the matrices $A^{*}$ and $B^{*}$ can be expressed as:
$[A^{*}]_{ik}$ $=\frac{\sum\limits_{(j,l)\in\mathcal{D}}\left[\mathbb{C}\right]_{(k,l),(i,j)}}{\sum\limits_{(j,l)\in\mathcal{D}}[B^{*}]_{lj}}\quad\text{and}\quad[B^{*}]_{lj}=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{m}\left[\mathbb{C}\right]_{(k,l),(i,j)}.$
Remark 2.7.
Notably, from Remark 2.5, the magnitude of $[B^{*}]_{lj}$ can be expressed as:
$\left\lvert[B^{*}]_{lj}\right\rvert=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{n}\left\lvert\left[\mathbb{C}\right]_{(k,l),(i,j)}\right\rvert.$
3 Analysis in the Noisy Setting
In this section we focus on the BMLR model. Given $T$ observations $(X_{t},Y_{t})_{t=1}^{T}$ that satisfy (1), our objective is to estimate the parameters $A^{*}\in\mathbb{R}^{n\times m}$ and $B^{*}\in\mathbb{R}^{q\times p}$.
From the observations $(X_{t},Y_{t})_{t=1}^{T}$, we construct the matrices $\mathbb{Y}\in\mathbb{R}^{T\times np}$ and $\mathbb{X}\in\mathbb{R}^{T\times mq}$ as follows:
$\mathbb{Y}:=\begin{pmatrix}\text{vec}(Y_{1})^{\top}\\ \text{vec}(Y_{2})^{\top}\\ \vdots\\ \text{vec}(Y_{T})^{\top}\end{pmatrix}\quad\text{and}\quad\mathbb{X}:=\begin{pmatrix}\text{vec}(X_{1})^{\top}\\ \text{vec}(X_{2})^{\top}\\ \vdots\\ \text{vec}(X_{T})^{\top}\end{pmatrix}.$
We assume that the design matrices $(X_{t})_{t=1}^{T}$ form a generating family of $\mathbb{R}^{m\times q}$. This assumption implies that $T\geq mq$ and ensures that $\mathbb{X}^{\top}\mathbb{X}\in\mathbb{R}^{mq\times mq}$ is invertible. Following Remark 2.2, this is a mild assumption.
We further define the unobserved noise matrix $\mathbb{E}\in\mathbb{R}^{T\times np}$ and the unobserved signal matrix $\mathbb{M}\in\mathbb{R}^{T\times np}$ as follows:
$\mathbb{E}:=\begin{pmatrix}\text{vec}(E_{1})^{\top}\\ \text{vec}(E_{2})^{\top}\\ \vdots\\ \text{vec}(E_{T})^{\top}\end{pmatrix}\quad\text{and}\quad\mathbb{M}:=\begin{pmatrix}\text{vec}(M_{1})^{\top}\\ \text{vec}(M_{2})^{\top}\\ \vdots\\ \text{vec}(M_{T})^{\top}\end{pmatrix}$ (5)
where $(E_{t})_{t=1}^{T}$ are the noise matrices defined in (1) and for $t\in[T]$, $M_{t}:=A^{*}X_{t}B^{*}.$
Using the structure of $\mathbb{X}\in\mathbb{R}^{T\times mq}$ we derive the unobserved matrix $\mathbb{C}\in\mathbb{R}^{mq\times np}$ as $\mathbb{C}:=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}.$ The relationship between $\mathbb{Y}$, $\mathbb{X}$, and the parameters $A^{*}$ and $B^{*}$ can now be expressed as:
$\mathbb{Y}=\mathbb{M}+\mathbb{E}=\mathbb{X}\mathbb{C}+\mathbb{E}.$ (6)
To analyze the influence of noise in the estimation process, we define $\mathbb{D}\in\mathbb{R}^{mq\times np}$ as $\mathbb{D}:=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{E}.$
Finally, we derive from the observations ^$\widehat{\mathbb{C}}\in\mathbb{R}^{mq\times np}$ as:
^$\widehat{\mathbb{C}}:=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{Y}.$ (7)
Using this definition, (6) leads to ^$\widehat{\mathbb{C}}=\mathbb{C}+\mathbb{D}.$ We assume that for all $(i,j,k)\in[n]\times[p]\times[q]$, the row sums $\sum_{k=1}^{m}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}$ are nonzero. This assumption ensures that the plug-in estimator for $A^{*}$ is well-defined. Notably, this assumption holds almost surely when the noise matrices $(E_{t})_{t=1}^{T}$ are drawn from a continuous distribution.
Leveraging the results from Corollary 2.6, we define the plug-in estimator $\hat{B}\in\mathbb{R}^{q\times p}$ of $B^{*}$, defined for all $(j,l)\in[p]\times[q]$ as:
$[\hat{B}]_{lj}=\dfrac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{k=1}^{m}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}.$ (8)
Similarly we define the plug-in estimator $\tilde{A}\in\mathbb{R}^{n\times m}$ of $A^{*}$, defined for all $(i,k)\in[n]\times[m]$, as follows:
$[\tilde{A}]_{ik}:=\dfrac{\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}}{\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[\tilde{B}]_{lj}^{(i)}}\quad\text{where}\quad[\tilde{B}]_{lj}^{(i)}:=\frac{1}{n-1}\sum\limits_{\begin{subarray}{c}r=1\\ r\neq i\end{subarray}}^{n}\sum\limits_{s=1}^{m}\left[\widehat{\mathbb{C}}\right]_{(s,l),(r,j)}.$ (9)
In the definition of $[\tilde{A}]_{ik}$, we ensure that the terms in the numerator do not appear in the denominator to preserve statistical independence of both terms. It is important to note that the entries of $\tilde{A}$ are defined as the ratio of random variables. While the behavior of such ratios has been studied in the literature, particularly in Gaussian cases (Geary,, 1930; Marsaglia,, 1965), the results obtained through this approach would require heavy assumptions and remain challenging to interpret in our context.
When the Gaussian variables in the ratio are centered, the distribution of the ratio is known as the Cauchy distribution (Johnson et al.,, 1995). However, for non-centered Gaussian variables, the probability density function of the ratio takes a significantly more complex form (Hinkley,, 1969), making the analysis cumbersome. Under certain conditions, it is possible to approximate the ratio with a normal distribution (Díaz-Francés & Rubio,, 2013), but this requires additional assumptions on the model and would still yield results that are difficult to decipher. Consequently, we opt for a different approach that avoids these complications while retaining interpretability.
Specifically, we observe that the plug-in estimator $\tilde{A}$ does not fully exploit a key property of the model: the entries of the matrix $A^{*}$ are constrained to lie between 00 and $1$. This additional structure could be leveraged to improve the estimator’s performance. Hence we define the estimator $\hat{A}\in\mathbb{R}^{n\times m}$ of $A^{*}$, defined for all $(i,k)\in[n]\times[m]$, as:
$[\hat{A}]_{ik}:=\max\left(0,\;\min\left([\tilde{A}]_{ik},1\right)\right).$ (10)
3.1 Dense Gaussian Case with Known Variance
First, we study the matrix normal case with known fixed variance under the ORT assumption. The matrix normal distribution generalizes the multivariate normal distribution to matrix-valued random variables (Gupta & Nagar,, 2018; Barratt,, 2018; Mathai et al.,, 2022).
Lemma 3.1.
Under the assumption that $\mathbb{X}^{\top}\mathbb{X}$ is full rank, the matrix ^$\widehat{\mathbb{C}}\in\mathbb{R}^{mq\times np}$ defined in (7) satisfies
^$\widehat{\mathbb{C}}\sim\mathcal{MN}_{mq\times np}\left(\mathbb{C},\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1},\Sigma_{c}\otimes\Sigma_{r}\right).$
Following the vast literature on linear regression and Gaussian sequence models (Tsybakov,, 2009; Giraud,, 2021; Rigollet & Hütter,, 2023; Chhor et al.,, 2024), we make the ORT assumption to capture a better understanding of the phenomenon at play.
Assumption 3.2 (ORT assumption).
We assume that the design matrix $\mathbb{X}$ satisfies $\dfrac{\mathbb{X}^{\top}\mathbb{X}}{T}=I_{mq}$.
We also assume that the covariance matrices are diagonal.
Assumption 3.3 (Diagonal Covariance matrices assumption).
We assume that the row-covariance matrix $\Sigma_{r}$ and the column-covariance matrix $\Sigma_{c}$ are diagonal. Namely we assume that there exist two positive numbers $\sigma_{r}^{2}$ and $\sigma_{c}^{2}$ such that $\Sigma_{r}=\sigma_{r}^{2}I_{n}$ and $\Sigma_{c}=\sigma_{c}^{2}I_{p}$.
We underline that under assumptions 3.2 and 3.3, Lemma 3.1 ensures that the entries of $\widehat{\mathbb{C}}$ are independent and normally distributed. We also note that Assumption 3.3 ensures that
$\Sigma_{c}\otimes\Sigma_{r}=\sigma^{2}I_{np},$
where $\sigma:=\sigma_{c}\sigma_{r}$. The following theorem establishes non-asymptotic upper bounds on the convergence rates of $\hat{B}$ under various norms.
Theorem 3.4.
Under Assumptions 3.2 and 3.3, the estimator $\hat{B}$ introduced in (8) satisfies the following non-asymptotic bounds for any $\epsilon>0$:
$\mathbb{P}\left[\|\hat{B}-B^{*}\|_{\max}>\epsilon\right]\leq$ $\frac{pq\sigma}{\epsilon}\sqrt{\frac{2m}{Tn\pi}}\exp\left(-\frac{Tn\epsilon^{2}}{2m\sigma^{2}}\right),$
$\mathbb{P}\left[\|\hat{B}-B^{*}\|_{\mathrm{op}}>\epsilon\right]\leq$ $2\exp\left(-\left(\frac{\epsilon\sqrt{nT}}{2\sigma\sqrt{m}}-\dfrac{\sqrt{p}+\sqrt{q}}{2}\right)^{2}\right),$
$\mathbb{P}\left[\|\hat{B}-B^{*}\|_{F}^{2}>\epsilon\right]\leq$ $2\exp\left(-\frac{Tn\epsilon}{3m\sigma^{2}}+\frac{2pq}{3}\right).$
Here $\|\cdot\|_{\max}$ is the elementwise maximum norm, $\|\cdot\|_{\mathrm{op}}$ is the operator norm, and $\|\cdot\|_{F}$ is the Frobenius norm.
We observe that the convergence rate of the estimator $\hat{B}$ exhibits the anticipated dependence on the sample size $T$, improving as the number of observations increases. Notably, our analysis reveals a ”blessing of dimensionality” effect, wherein the convergence rate accelerates as the row dimension $n$ of the observed target matrices $(Y_{t})_{t=1}^{T}$ grows.
Conversely, the convergence rate is negatively affected by the size $pq$ of $B^{*}$. Furthermore, the variance parameter $\sigma:=\sigma_{c}\sigma_{r}$ also exerts a detrimental influence on the convergence rate, as intuitively expected.
The following Lemma provides a probabilistic control over the event where the plug-in estimator $\tilde{A}$, defined in (9), coincides with its modified version $\hat{A}$ defined in (10). To avoid sparsity in the parameter $A^{*}$ we first assume that its entries are strictly bounded from above by $1$. We also assume that the sum of the entries of $B^{*}$ is positive.
Assumption 3.5.
We assume that for all $(i,k)\in[n]\times[m]$ we have $10<[A^{*}]_{ik}<$ 1. In addition we assume that $\beta^{*}>0$ where $\beta^{*}:=\dfrac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[B^{*}\right]_{lj}$.
We now state the corresponding Lemma.
Lemma 3.6.
Under Assumptions 3.2, 3.3 and 3.5, the estimators $\tilde{A}$ and $\hat{A}$ introduced in (9) and (10) satisfy the following property for any $(i,k)\in[n]\times[m]$:
$\mathbb{P}\left[[\tilde{A}]_{ik}\neq[\hat{A}]_{ik}\right]\leq$ $\frac{\tilde{\sigma}}{\mu_{ik}\sqrt{2pqT\pi}}\exp\left(-\frac{Tpq\mu_{ik}^{2}}{2\sigma_{1}^{2}}\right)+\frac{\sigma}{\nu_{ik}\sqrt{2pqT\pi}}\exp\left(-\frac{Tpq\nu_{ik}^{2}}{2\sigma^{2}}\right),$
where $\mu_{ik}:=\frac{1}{pq}\sum_{l=1}^{q}\sum_{j=1}^{p}[B^{*}]_{lj}\left(1-[A^{*}]_{ik}\right)$, $\nu_{ik}:=[A^{*}]_{ik}\beta^{*}$ and $\tilde{\sigma}:=\sigma\sqrt{1+\frac{m}{n-1}}.$
We first note that under Assumption 3.5, the quantities $\mu_{ik}$ and $\nu_{ik}$ are positive. We then observe that as the sample size $T$ increases, the plug-in estimator $\tilde{A}$ is more likely to coincide with its modified version $\hat{A}$. The intuition behind this result is straightforward: as the sample size grows, $[\tilde{A}]_{ik}$ converges to $[A^{*}]_{ik}$, which inherently lies between 0 and 1. Consequently, the modifications introduced in $\hat{A}$ become unnecessary as $[\tilde{A}]_{ik}$ naturally satisfies the constraints of the model. A similar phenomenon occurs as $p$, the number of columns of the response matrices, and $q$, the number of columns of the predictor matrices, increase. Larger $p$ and $q$ effectively provide more information about the structure of the model, leading to improved accuracy of the plug-in estimator $\tilde{A}$ and reducing the need for corrections by $\hat{A}$.
The following theorem provides a finite-sample analysis of the performance of $\hat{A}$.
Theorem 3.7.
Under Assumptions 3.2, 3.3 and 3.5, the estimator $\hat{A}$ introduced in (10) satisfies for any $\epsilon>0$:
$\mathbb{P}\left[\left\|\hat{A}-A^{*}\right\|_{\max}>\frac{2\epsilon}{\left\lvert\beta^{*}\right\rvert}\right]$ $\leq\frac{nm\sigma\sqrt{2}}{\epsilon\sqrt{pqT\pi}}\left[\exp\left(-\frac{Tpq\epsilon^{2}}{2\sigma^{2}}\right)+\sqrt{\frac{m}{n-1}}\exp\left(-\frac{Tpq(n-1)\epsilon^{2}}{2m\sigma^{2}}\right)\right]$
$+\frac{nm}{\sqrt{2pqT\pi}}\left[\frac{\tilde{\sigma}\exp\left(-\dfrac{Tpq\mu^{2}}{2\sigma_{1}^{2}}\right)}{\mu}+\frac{\sigma\exp\left(-\dfrac{Tpq\nu^{2}}{2\sigma^{2}}\right)}{\nu}\right]$
where $\nu:=\min_{i,k}\nu_{ik}$, $\mu:=\min_{i,k}\mu_{ik}$ with $\mu_{ik}$, $\nu_{ik}$ and $\tilde{\sigma}$ defined in Lemma 3.6.
We observe that the finite-sample performance of the estimator $\hat{A}$ improves with the sample size $T$, reflecting the benefit of more observations. Additionally, the convergence rate is positively influenced by increases in the column dimension $p$ of the observed target matrices $(Y_{t})_{t=1}^{T}$ and the column dimension $q$ of the predictor matrices $(X_{t})_{t=1}^{T}$. This reflects a ”blessing of dimensionality” effect, as additional columns provide richer information for the estimation process. Notably, this behavior contrasts with that of $\hat{B}$, as detailed in Theorem 3.4, where increases in $p$ and $q$ have a detrimental impact. Moreover, the magnitude of $\beta^{*}$ plays a role in the convergence rate, with larger values of $|\beta^{*}|$ leading to faster convergence.
Conversely, the performance of $\hat{A}$ deteriorates as the size $nm$ of $A^{*}$ increases. Finally, the variance parameter $\sigma:=\sigma_{r}\sigma_{c}$ negatively affect the convergence rate, higher noise levels degrading the accuracy of the estimator. These observations underline the nuanced interplay between dimensionality and noise in determining the performance of $\hat{A}$.
3.2 Sparse Gaussian Case with Known Variance
In the second part, we extend our analysis to incorporate sparsity assumptions on the parameters $A^{*}$ and $B^{*}$. By leveraging the sparse structure, we aim to develop estimation strategies tailored for high-dimensional settings, where many entries of the parameters are expected to be zero or near-zero. This sparse framework addresses practical scenarios where dimensionality reduction and interpretability are critical.
Building on ideas developed in (Giraud,, 2021; Rigollet & Hütter,, 2023) for vector estimation, we propose the hard-thresholding estimator $\hat{B}^{S}$, defined as follows for all $(l,j)\in[q]\times[p]$:
$[\hat{B}^{S}]_{lj}=\begin{cases}[\hat{B}]_{lj},&\text{if }\lvert[\hat{B}]_{lj}\rvert>2\tau,\\ 0,&\text{if }\lvert[\hat{B}]_{lj}\rvert\leq 2\tau,\end{cases}$ (11)
where $[\hat{B}]_{lj}$ are the entries of the initial estimator $\hat{B}$ defined in (8), and $\tau>0$ is a user-defined threshold. The hard-thresholding operation enforces sparsity by setting small entries of $\hat{B}$ to zero, aligning the estimator with the sparse structure of $B^{*}$.
The following theorem establishes a non-asymptotic upper bound on the convergence rate of $\hat{B}^{S}$ for the Frobenius norm under sparsity assumptions on the parameter $B^{*}$.
Theorem 3.8.
Under Assumptions 3.2 and 3.3, for any $\delta\in(0,1)$, the estimator $\hat{B}^{S}$ introduced in (11), with the threshold:
$\tau:=\sigma\sqrt{\frac{2m}{nT}}\left(\sqrt{\log(2pq)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right),$
enjoys the following non-asymptotic properties on the same event holding with probability at least $1-\delta$:
  1. 1.
    If $\|B^{*}\|_{0}$ denotes the number of nonzero coefficients in $B^{*}$, then:
    $\|\hat{B}^{S}-B^{*}\|_{F}^{2}<16\left\|B^{*}\right\|_{0}\tau^{2}.$
  2. 2.
    If the minimum magnitude of the entries of $B^{*}$ satisfies:
    $\min_{l\in[q],j\in[p]}\lvert[B^{*}]_{lj}\rvert>3\tau,$
    then the support of $\hat{B}^{S}$ perfectly matches that of $B^{*}$:
    $\text{supp}(\hat{B}^{S})=\text{supp}(B^{*}),$
    where $\text{supp}(\cdot)$ denotes the set of indices corresponding to the nonzero entries of a matrix.
Theorem 3.8 highlights the performance of the sparse estimator $\hat{B}^{S}$ under sparsity assumptions on the true parameter $B^{*}$. The convergence rate of $\hat{B}^{S}$ improves as the sparsity of $B^{*}$ increases, demonstrating the benefits of leveraging sparse structures. The threshold $\tau$ exhibits favorable scaling with the sample size $T$ and row dimension $n$, both of which contribute to reducing $\tau$, enhancing the estimator’s performance. Conversely, $\tau$ increases with the dimensions $p$ and $q$, reflecting the greater challenge of estimation in higher-dimensional settings. Moreover the threshold $\tau$ scales with $\sigma$, capturing the adverse impact of higher noise levels on the estimation accuracy. This emphasizes the observations from Theorem 3.4. Finally the condition on the minimum magnitude of the entries of $B^{*}$ ensures that the threshold $\tau$ enables exact recovery of the support of $B^{*}$, with high probability.
We now propose the hard-thresholding estimator $\hat{A}^{S}$, defined as follows for all $(i,k)\in[n]\times[m]$:
$[\hat{A}^{S}]_{ik}=\begin{cases}[\hat{A}]_{ik},&\text{if }\lvert\hat{\gamma}_{ik}\rvert>2\tau,\\ 0,&\text{if }\lvert\hat{\gamma}_{ik}\rvert\leq 2\tau,\end{cases}$ (12)
where:
$\hat{\gamma}_{ik}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}.$
The following theorem establishes a non-asymptotic upper bound on the convergence rate of $\hat{A}^{S}$ for the Frobenius norm under sparsity assumptions on the parameter $A^{*}$. We recall that the quantity $\beta^{*}\in\mathbb{R}$ is defined as $\beta^{*}:=\dfrac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[B^{*}\right]_{lj}$.
Theorem 3.9.
Under Assumptions 3.2 and 3.3, for any $\delta\in(0,1)$, the estimator $\hat{A}^{S}$ defined in (12), with the threshold:
$\tau:=\sigma\sqrt{\frac{2}{Tpq}}\left(\sqrt{\log(2nm)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right),$
satisfies the following non-asymptotic properties on an event holding with probability at least $1-2\delta$:
  1. 1.
    If $\|A^{*}\|_{0}$ denotes the number of nonzero coefficients in $A^{*}$, then:
    $\|\hat{A}^{S}-A^{*}\|_{F}^{2}\leq\lvert\beta^{*}\rvert^{-2}\|A^{*}\|_{0}\left(2t_{\delta}+3\tau\right)^{2},$
    where $t_{\delta}$ satisfies:
    $\frac{\sigma\sqrt{2m}}{t_{\delta}\sqrt{npqT\pi}}\exp\left(-\frac{Tpqt_{\delta}^{2}}{2m\sigma^{2}}\right)\quad+\frac{\sigma\sqrt{2}}{t_{\delta}\sqrt{pqT\pi}}\exp\left(-\frac{Tpqt_{\delta}^{2}}{2\sigma^{2}}\right)=\delta.$
  2. 2.
    If the minimum magnitude of the entries of $A^{*}$ satisfies:
    $\min_{i\in[n],k\in[m]}\lvert[A^{*}]_{ik}\rvert>3\tau,$
    and if $13\beta^{*}>$ 1, then the support of $\hat{A}^{S}$ perfectly matches that of $A^{*}$:
    $\text{supp}(\hat{A}^{S})=\text{supp}(A^{*}).$
Remark 3.10.
The parameter $t_{\delta}$, which determines the concentration properties of the estimator, decreases as $T$, $p$, and $q$ increase. Consequently, the estimator $\hat{A}^{S}$ benefits from improved performance as the sample size $T$ and latent dimensions $p$ and $q$ grow.
Theorem 3.9 characterizes the non-asymptotic properties of the sparse estimator $\hat{A}^{S}$ under sparsity assumptions on the true parameter matrix $A^{*}$.
The first result provides a Frobenius norm error bound that scales with the sparsity level $\|A^{*}\|_{0}$. This bound is inversely related to the squared magnitude of $\beta^{*}$ (the average entry in $B^{*}$), indicating that stronger signals in the underlying parameter matrix $B^{*}$ lead to improved estimation accuracy. This is similar to the phenomenon described in Theorem 3.7. The error bound also depends on both the threshold parameter $\tau$ and the concentration parameter $t_{\delta}$, which capture the impact of noise and sample size on the estimation performance. The threshold $\tau$ exhibits several important dependencies. It decreases with the sample size $T$, reflecting improved estimation with more observations. It similarly decreases with dimensions $p$ and $q$, showcasing a beneficial effect of higher dimensionality. Conversely it scales with the noise level $\sigma$, capturing the detrimental impact of increased noise levels. Finally it grows logarithmically with the matrix dimensions $n$ and $m$, indicating a mild sensitivity to the size of $A^{*}$.
The second result establishes conditions for perfect support recovery of $A^{*}$. Specifically, when the minimum magnitude of the nonzero entries in $A^{*}$ exceeds three times the threshold $\tau$, and the average effect $\beta^{*}$ is sufficiently large ($13\beta^{*}>$ 1), the sparse estimator exactly recovers the support of $A^{*}$.
As noted in Remark 3.10, the concentration parameter $t_{\delta}$ decreases with larger values of $T$, $p$, and $q$. This property, combined with the similar behavior of $\tau$, demonstrates that the sparse estimator benefits from both increased sample size and higher latent dimensions, a particularly favorable characteristic for high-dimensional settings.
4 Numerical Simulations
The code used to perform all numerical simulations and generate the results presented in this paper is publicly available on GitHub: https://github.com/nayelbettache/BMLR .
4.1 Synthetic data
In this section, we evaluate the performance of the proposed estimators $\hat{A}$ and $\hat{B}$ through numerical simulations. The simulations are conducted using a Python implementation, and the following setup is used to generate the required matrices and data.
Simulation Setup
The simulations involve the generation of matrices $A^{*}$, $B^{*}$, $X_{t}$, $E_{t}$ and $Y_{t}$. By default, the parameters are set as $n=15$, $m=13$, $p=14$, $q=12$, $T=1000$, and $\sigma=1$. These default parameter values are adjusted to analyze the effects of $n$, $m$, $p$, $q$, $T$, and $\sigma$ on the performance of the proposed estimators. $A^{*}$ is a $n\times m$ matrix with random entries sampled from a uniform distribution over $[0,1)$ and rows then normalized to sum to 1 (L1-normalized). $B^{*}$ is a $q\times p$ matrix with random entries sampled from a uniform distribution over $[0,1)$. $(X_{t})$ is a sequence of $T$ matrices of size $m\times q$ with random entries sampled from a uniform distribution over $[0,1)$. $(E_{t})$ is a sequence of $T$ noise matrices of size $n\times p$, with entries drawn from a Gaussian distribution with mean 0 and standard deviation $\sigma=1$. $(Y_{t})$ is a sequence of $T$ observation matrices, where $Y_{t}=A^{*}X_{t}B^{*}+E_{t}$.
Estimation and Evaluation
The estimators $\hat{A}$ and $\hat{B}$ are computed using (10) and (8) respectively. To evaluate the performance of the estimators we vary the parameters $n$, $m$, $p$, $q$, and $T$ to observe their impact on the estimation accuracy. For each parameter setting, we compute the squared Frobenius norm of the errors:
$\|\hat{A}-A\|_{F}^{2}\quad\text{and}\quad\|\hat{B}-B\|_{F}^{2}.$
These errors are plotted as functions of the varying parameters in Figures 1 and 2, respectively.
Validation of Theoretical Properties
From the plots, we observe that as $T$ increases, the errors in $\hat{A}$, Figure 1(c) and $\hat{B}$, Figure 2(c), decrease, indicating improved estimation accuracy with more data. As $p$ and $q$ increase, the errors in $\hat{A}$ decrease, Figures 1(d) and 1(e) and the errors in $\hat{B}$ increase, Figures 2(d) and 2(e). As $m$ increases, both the errors in $\hat{A}$ and $\hat{B}$ increase, Figures 1(b) and 2(b) respectively. As $n$ increases, the errors in $\hat{A}$ increase, Figure 1(a) and the errors in $\hat{B}$ decrease, Figure 2(a).
These results confirm the theoretical properties of the estimators, detailed in Theorems 3.7 and 3.4 respectively.
Additionally, we verify the results stated in Corollary 2.6 by comparing them with numerical simulations. While exact equality cannot be assessed due to inherent numerical precision limitations, the computed values closely match, with differences within $10^{-8}$. This consistency provides strong support for the validity of our theoretical claims.
(a) EV w.r.t $n$
(b) EV w.r.t $m$
(c) EV w.r.t $T$
(d) EV w.r.t $p$
(e) EV w.r.t $q$
Figure 1: Error variation (EV) of the Frobenius norm errors of $\hat{A}$ for estimating $A^{*}$ with respect to (w.r.t.) different parameters
(a) EV w.r.t $n$
(b) EV w.r.t to $m$
(c) EV w.r.t to $T$
(d) EV w.r.t to $p$
(e) EV w.r.t to $q$
Figure 2: Error variation (EV) of the Frobenius Norm Errors of $\hat{B}$ for estimating $B^{*}$ with respect to (w.r.t.) different parameters
4.2 Real-world data
In addition to the synthetic simulations, we evaluate the performance of our proposed methods using real-world data. For this purpose, we use the CIFAR-10 dataset, a collection of natural images widely used in machine learning and computer vision. The goal is to apply our methodology to simulate noisy transformations of the images and assess the effectiveness of the proposed correction techniques.
The CIFAR-10 dataset consists of 50,000 training images and 10,000 test images, where each image is an RGB image of size $32\times 32\times 3$, representing three color channels: red, green, and blue. The pixel values are normalized to the range $[0,1]$ to ensure consistency in computations.
Noisy Transformations
We simulate noisy transformations of the images using left and right multiplications with matrices $A^{*}$ and $B^{*}$. These matrices introduce controlled noise into the images while maintaining specific structural properties. Specifically, matrix $A^{*}$ is derived from $A=I_{32}+\epsilon E_{1}$, where $E_{1}$ is a $32\times 32$ matrix with independent and identically distributed entries sampled from a standard Gaussian distribution. The matrix $A^{*}$ is obtained by normalizing each row of $A$ to sum to 1. Similarly, matrix $B^{*}$ is given by $B^{*}=I_{32}+\epsilon E_{2}$, where $E_{2}$ is a $32\times 32$ matrix with i.i.d. entries sampled from a standard Gaussian distribution. The parameter $\epsilon$ controls the noise level in both transformations.
For both training and test images, the noisy transformation is applied independently to each color channel. The transformation for a given channel $c\in\{1,2,3\}$ (corresponding to red, green, and blue) is defined as:
$X_{\text{noisy}}^{(c)}=(A^{*})^{-1}\cdot X_{\text{or}}^{(c)}\cdot(B^{*})^{-1},$
where $X_{\text{or}}^{(c)}$ represents the $c^{\text{th}}$ channel of the original image and $X_{\text{noisy}}^{(c)}$ is the corresponding noisy version.
Correction Process
The correction process is learnt using only the training set. From the noisy transformed images from the training set, we estimate the correction matrices $\hat{A}_{\text{train}}^{(c)}$ and $\hat{B}_{\text{train}}^{(c)}$ using Equations (10) and (8) respectively. Each color channel is processed independently to obtain these estimates.
Once the correction matrices $\hat{A}_{\text{train}}^{(c)}$ and $\hat{B}_{\text{train}}^{(c)}$ are computed, they are applied to the noisy test images to reconstruct the corrected test images. For each channel in a test image, the corrected channel is computed as:
$X_{\text{corr}}^{(c)}=\hat{A}_{\text{train}}^{(c)}\cdot X_{\text{noisy}}^{(c)}\cdot\hat{B}_{\text{train}}^{(c)}.$
Figure 3 illustrates an example from the test set, showcasing three versions of the same image: the original image (left), its noisy counterpart generated with a noise factor of $\epsilon=0.02$ (middle), and the corresponding corrected image obtained after applying the proposed denoising process (right). The noisy image clearly exhibits significant distortion due to the added noise, while the corrected image demonstrates the effectiveness of the estimation procedure in restoring the original image quality.
Figure 3: Example of original, noisy, and corrected versions of the $11^{\text{th}}$ image from the test set for $\epsilon=0.02$.
Evaluation of Correction Quality
To evaluate the effectiveness of the correction process, we compute the Frobenius norm distances between the original image and its noisy or corrected versions:
$\text{D}_{\text{o,n}}=\sum\limits_{c=1}^{3}\|X_{\text{or}}^{(c)}-X_{\text{noisy}}^{(c)}\|_{F}^{2},\quad\text{D}_{\text{o,c}}=\sum\limits_{c=1}^{3}\|X_{\text{or}}^{(c)}-X_{\text{corr}}^{(c)}\|_{F}^{2}.$
For the test set example used in Figure 3, the computed distances are: $\text{D}_{\text{o,n}}=215.44$ and $\text{D}_{\text{o,c}}=79.88$.
To provide a broader perspective, we plot the distance metrics $\text{D}_{\text{o,n}}$ and $\text{D}_{\text{o,c}}$ as functions of the noise factor $\epsilon$, averaged over the entire test set. The results are shown in Figure 4.
Figure 4: $\text{D}_{\text{o,n}}$ in blue and $\text{D}_{\text{o,c}}$ in orange, averaged on the test set, as functions of the noise factor $\epsilon$. Error bars indicate standard deviations.
Conclusion
The results demonstrate that the proposed correction process effectively mitigates the impact of noise. The corrected images closely approximate the original images, as shown by both qualitative (Figure 3) and quantitative (Figure 4) metrics. This methodology generalizes well to real-world data, underscoring the applicability of our framework beyond synthetic simulations.
Appendix A Proofs
A.1 Proofs of Section 2
Proof of Lemma 2.1.
From the noiseless bivariate matrix-valued linear regression (NBMLR) model (3), we have
$M_{(k,l)}=A^{*}E_{(k,l)}B^{*},$
where $E_{(k,l)}$ is the matrix with a single entry of $1$ at the $(k,l)$-th position and zeros elsewhere. Expanding this relationship element-wise gives:
$[M_{(k,l)}]_{ij}=[A^{*}]_{ik}\cdot[B^{*}]_{lj},\;\forall(i,j,k,l)\in[n]\times[p]\times[m]\times[q].$
Step 1: Recovery of $[B^{*}]_{lj}$.
Summing over $k\in[m]$ for fixed $l,j,i$, we observe that:
$\sum_{k=1}^{m}[M_{(k,l)}]_{ij}=\sum_{k=1}^{m}[A^{*}]_{ik}\cdot[B^{*}]_{lj}.$
Since the rows of $A^{*}$ have unit $L_{1}$-norm, it follows that:
$[B^{*}]_{lj}=\sum_{k=1}^{m}[M_{(k,l)}]_{ij},\;\forall(l,j,i)\in[q]\times[p]\times[n].$
Step 2: Recovery of $[A^{*}]_{ik}$.
From the model equation, for each $(k,l)$, we isolate $[A^{*}]_{ik}$ by dividing $[M_{(k,l)}]_{ij}$ by $[B^{*}]_{lj}$, provided that $[B^{*}]_{lj}\neq 0$:
$[A^{*}]_{ik}=\frac{[M_{(k,l)}]_{ij}}{[B^{*}]_{lj}},\;\forall(i,k)\in[n]\times[m],\;\text{where}\;[B^{*}]_{lj}\neq 0.$
Proof of Lemma 2.3.
The goal is to show that the matrix $\mathbb{C}$ can be expressed as:
$\mathbb{C}=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}.$
This is equivalent to demonstrating that, for all $(i,j)\in[n]\times[p]$, the following equality holds:
$[A^{*}X_{t}B^{*}]_{ij}=\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{q}[X_{t}]_{kl}\left[\mathbb{C}\right]_{(k,l),(i,j)}.$ (13)
By construction, the entries of $\mathbb{C}$ are given as:
$\left[\mathbb{C}\right]_{(k,l),(i,j)}=[A^{*}]_{ik}\cdot[B^{*}]_{lj}.$
Substituting this into the right-hand side of (13), we have:
$\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{q}[X_{t}]_{kl}\left[\mathbb{C}\right]_{(k,l),(i,j)}=\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{q}[X_{t}]_{kl}\cdot[A^{*}]_{ik}\cdot[B^{*}]_{lj}.$
On the left-hand side of (13), by definition of the matrix product, we know:
$[A^{*}X_{t}B^{*}]_{ij}=\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{q}[A^{*}]_{ik}\cdot[X_{t}]_{kl}\cdot[B^{*}]_{lj}.$
Since the two expressions above are identical, we conclude that:
$[A^{*}X_{t}B^{*}]_{ij}=\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{q}[X_{t}]_{kl}\left[\mathbb{C}\right]_{(k,l),(i,j)}.$
This establishes that $\mathbb{M}=\mathbb{X}\mathbb{C}$. As $\mathbb{X}$ spans $\mathbb{R}^{mq}$, the equality $\mathbb{C}=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}$ follows directly. ∎
Proof of Proposition 2.4.
From Lemma 2.3, the matrix $\mathbb{C}$ can be expressed as:
$\mathbb{C}=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}.$
This means that $\mathbb{C}$ can be computed directly from the observations $(X_{t},M_{t})_{t\in[T]}$, provided that $\mathbb{X}^{\top}\mathbb{X}$ is invertible, which is ensured by the assumption that the design matrices $(X_{t})_{t=1}^{T}$ form a generating family of $\mathbb{R}^{m\times q}$.
Step 1: Relating $\mathbb{C}$ to $A^{*}$ and $B^{*}$.
By the definition of $\mathbb{C}$, each entry $\left[\mathbb{C}\right]_{(k,l),(i,j)}$ corresponds to:
$\left[\mathbb{C}\right]_{(k,l),(i,j)}=[A^{*}]_{ik}\cdot[B^{*}]_{lj},$
for all $i\in[n]$, $k\in[m]$, $l\in[q]$, and $j\in[p]$. This bilinear structure allows us to recover $A^{*}$ and $B^{*}$ separately by exploiting their roles in this product.
Step 2: Recovering $B^{*}$.
To isolate $[B^{*}]_{lj}$, we sum $\left[\mathbb{C}\right]_{(k,l),(i,j)}$ over all $k\in[m]$ for a fixed $l,j,i$. Specifically:
$\sum_{k=1}^{m}\left[\mathbb{C}\right]_{(k,l),(i,j)}=\sum_{k=1}^{m}\left([A^{*}]_{ik}\cdot[B^{*}]_{lj}\right).$
Since the rows of $A^{*}$ satisfy the $L_{1}$-norm constraint (i.e., $\sum_{k=1}^{m}[A^{*}]_{ik}=1$), the summation simplifies to:
$\sum_{k=1}^{m}\left[\mathbb{C}\right]_{(k,l),(i,j)}=[B^{*}]_{lj}.$
This establishes the second equation in the proposition:
$[B^{*}]_{lj}=\sum_{k=1}^{m}\left[\mathbb{C}\right]_{(k,l),(i,j)}.$
Step 3: Recovering $A^{*}$.
To isolate $[A^{*}]_{ik}$, we use the bilinear relationship:
$\left[\mathbb{C}\right]_{(k,l),(i,j)}=[A^{*}]_{ik}\cdot[B^{*}]_{lj}.$
For a fixed $i,k,l,j$, we can solve for $[A^{*}]_{ik}$ provided that $[B^{*}]_{lj}\neq 0$:
$[A^{*}]_{ik}=\frac{\left[\mathbb{C}\right]_{(k,l),(i,j)}}{[B^{*}]_{lj}}.$
This establishes the first equation in the proposition:
$[A^{*}]_{ik}=\frac{\left[\mathbb{C}\right]_{(k,l),(i,j)}}{[B^{*}]_{lj}},\quad\text{for}\quad[B^{*}]_{lj}\neq 0.$
Proof of Corollary 2.6.
The result on $B^{*}$ follows immediately from Proposition 2.4. To prove the result on $A^{*}$, we need to prove that for any $n\in\mathbb{N}^{*}$, if there exist $(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{R}^{n}$ and $(\beta_{1},\ldots,\beta_{n})\in\mathbb{R}^{n}$ such that:
$\gamma_{n}:=\frac{\alpha_{1}}{\beta_{1}}=\cdots=\frac{\alpha_{n}}{\beta_{n}},$
then it follows that:
$\gamma_{n}=\frac{\sum\limits_{i=1}^{n}\alpha_{i}}{\sum\limits_{j=1}^{n}\beta_{j}}.$
Specifically, since the entries of $A^{*}$ can be expressed as different ratios, all being equal, applying this result to the equations satisfied by $A^{*}$ in Proposition 2.4 completes the proof.
We prove this result by induction on $n$.
Initially at $n=1$ the result is trivially true. Assume the statement holds at step $n$, i.e., if
$\text{if}\quad\gamma_{n}:=\frac{\alpha_{1}}{\beta_{1}}=\cdots=\frac{\alpha_{n}}{\beta_{n}},\quad\text{then}\quad\gamma_{n}=\frac{\sum\limits_{i=1}^{n}\alpha_{i}}{\sum\limits_{j=1}^{n}\beta_{j}}.$
We now prove that the statement holds at step $n+1$.
Suppose:
$\gamma_{n+1}:=\frac{\alpha_{1}}{\beta_{1}}=\cdots=\frac{\alpha_{n+1}}{\beta_{n+1}}.$
By the definition of $\gamma_{n+1}$ and the assumption of step $n$ being true, we can write:
$\frac{\sum\limits_{i=1}^{n}\alpha_{i}}{\sum\limits_{j=1}^{n}\beta_{j}}=\gamma_{n+1}\quad\text{and}\quad\frac{\alpha_{n+1}}{\beta_{n+1}}=\gamma_{n+1}.$
This allows to write:
$\sum\limits_{i=1}^{n}\alpha_{i}=\left(\sum\limits_{j=1}^{n}\beta_{j}\right)\frac{\alpha_{n+1}}{\beta_{n+1}}.$
Hence we deduce:
$\frac{\sum\limits_{i=1}^{n}\alpha_{i}+\alpha_{n+1}}{\sum\limits_{j=1}^{n}\beta_{j}+\beta_{n+1}}=\frac{\left(\sum\limits_{j=1}^{n}\beta_{j}\right)\frac{\alpha_{n+1}}{\beta_{n+1}}+\alpha_{n+1}}{\sum\limits_{j=1}^{n}\beta_{j}+\beta_{n+1}}=\alpha_{n+1}\cdot\frac{\frac{\left(\sum\limits_{j=1}^{n}\beta_{j}\right)}{\beta_{n+1}}+1}{\sum\limits_{j=1}^{n}\beta_{j}+\beta_{n+1}}.$
Finally multiplying the numerator by $\beta_{n+1}$ and dividing $\alpha_{n+1}$ by the same factor provides:
$\frac{\sum\limits_{i=1}^{n}\alpha_{i}+\alpha_{n+1}}{\sum\limits_{j=1}^{n}\beta_{j}+\beta_{n+1}}=\frac{\alpha_{n+1}}{\beta_{n+1}}=\gamma_{n+1}.$
Thus, the statement holds at step $n+1$. ∎
A.2 Proofs of Section 3
Proof of Lemma 3.1.
We first recall that, by construction, we have:
^$\widehat{\mathbb{C}}=\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{M}+\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}\mathbb{E}=\mathbb{C}+\mathbb{D}.$
Next we prove the following, which will conclude:
$\mathbb{D}\sim\mathcal{MN}_{mq\times np}\left(0_{mq\times np},\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1},\Sigma_{c}\otimes\Sigma_{r}\right).$
First we recall that from model (1) the noise matrices $(E_{t})_{t=1}^{T}$ are independent and follow the same matrix normal distribution, namely $\mathcal{MN}_{n\times p}\left(0_{n\times p},\Sigma_{r},\Sigma_{c}\right)$. Hence by definition of the matrix normal distribution:
[…]
This leads to derive the distribution of the matrix $\mathbb{E}\in\mathbb{R}^{T\times np}$ defined in (5) as follows:
$\mathbb{E}\sim\mathcal{MN}_{T\times np}\left(0_{T\times np},I_{T\times T},\Sigma_{c}\otimes\Sigma_{r}\right).$
Finally, the matrix $\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}\mathbb{X}^{\top}$ being of full rank $mq$ and $\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1}$ being symmetric, by definition of the matrix $\mathbb{D}\in\mathbb{R}^{mq\times np}$ and by property of affine transformations of matrix gaussian distributions, we conclude that
$\mathbb{D}\sim\mathcal{MN}_{mq\times np}\left(0_{T\times np},\left(\mathbb{X}^{\top}\mathbb{X}\right)^{-1},\Sigma_{c}\otimes\Sigma_{r}\right).$
Proof of Theorem 3.4.
We start by deriving the finite-sampled inequalities on $\hat{B}$. From Lemma 3.1 and Assumption 3.2 we deduce that $\text{vec}(\widehat{\mathbb{C}})\sim\mathcal{N}_{mqnp}\left(\text{vec}(\mathbb{C}),\dfrac{\Sigma_{c}\otimes\Sigma_{r}}{T}\right)$. Assumption 3.3 then ensures that all entries of $\text{vec}(\widehat{\mathbb{C}})$ are independent and follow the same Gaussian distribution. Hence for all $(k,l,i,j)\in[m]\times[q]\times[n]\times[p]$ we have ^$\widehat{\mathbb{C}}_{(k,l),(i,j)}-\mathbb{C}_{(k,l),(i,j)}\overset{i.i.d}{\sim}\mathcal{N}\left(0,\dfrac{\sigma^{2}}{T}\right)$. Using the results from Proposition 2.4 and (8) lead to, for all $(l,j)\in[q]\times[p]$,
$[\hat{B}]_{lj}-[B^{*}]_{lj}\overset{i.i.d}{\sim}\mathcal{N}\left(0,\dfrac{m\sigma^{2}}{nT}\right).$ (14)
Mill’s inequality, Theorem C.1, ensures that for all $(l,j)\in[q]\times[p]$, for any $\epsilon>0$,
$\mathbb{P}\left[\left\lvert[\hat{B}]_{lj}-[B^{*}]_{lj}\right\rvert>\epsilon\right]\leq\dfrac{\sigma}{\epsilon}\sqrt{\dfrac{2m}{nT\pi}}\exp\left(-\dfrac{nT\epsilon^{2}}{2m\sigma^{2}}\right).$
The first inequality is then deduced by using a union bound.
The second inequality is immediately derived from the concentration of extreme singular values of Gaussian matrices with independent entries, see Corollary 5.35 in (Vershynin,, 2010).
For the third inequality, we first note that $\|\hat{B}-B^{*}\|_{F}^{2}=\sum\limits_{l=1}^{q}\sum\limits_{j=1}^{p}\left([\hat{B}]_{lj}-[B^{*}]_{lj}\right)^{2}$ follows a $\dfrac{m\sigma^{2}}{nT}\cdot\chi^{2}(pq)$ distribution. Then, Lemma 1 from (Laurent & Massart,, 2000) ensures that a random variable $Z$ following a $\chi^{2}(pq)$ distribution satisfies, for any $\epsilon>0$:
$\mathbb{P}\left(Z\geq pq+2\sqrt{pq\epsilon}+2\epsilon\right)\leq\exp(-\epsilon).$
Hence we deduce that
$\mathbb{P}\left(\|\hat{B}-B^{*}\|_{F}^{2}\geq\dfrac{m\sigma^{2}}{nT}\left(pq+2\sqrt{pq\epsilon}+2\epsilon\right)\right)\leq\exp(-\epsilon).$
We notice that, for any $\epsilon>0$, the following inequality holds:
$\dfrac{m\sigma^{2}}{nT}\left(pq+2\sqrt{pq\epsilon}+2\epsilon\right)\leq\dfrac{m\sigma^{2}}{nT}\left(2pq+3\epsilon\right)$
It leads to:
$\mathbb{P}\left(\|\hat{B}-B^{*}\|_{F}^{2}\geq\dfrac{m\sigma^{2}}{nT}\left(2pq+3\epsilon\right)\right)\leq\exp(-\epsilon).$
The stated result follows by a change of variable. ∎
Proof of Lemma 3.6.
Let us fix $(i,k)\in[n]\times[m]$. Using the definitions of $\hat{A}$ and $\tilde{A}$ from (10) and (9), respectively, we find that $[\tilde{A}]_{ik}=[\hat{A}]_{ik}$ if and only if $[\tilde{A}]_{ik}\in[0,1]$. Hence $[\tilde{A}]_{ik}=[\hat{A}]_{ik}$ if and only if the event $\mathcal{A}$ holds, where:
$\mathcal{A}:=\left\{0\leq\dfrac{\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}}{\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[\tilde{B}]_{lj}^{(i)}}\leq 1\right\},$
with $[\tilde{B}]_{lj}^{(i)}$ defined in (9) This event is realized if and only if the event $\mathcal{A}_{+}$ or the event $\mathcal{A}_{-}$ holds, where:
$\mathcal{A}_{+}:=\left\{0\leq\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}\leq\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[\tilde{B}]_{lj}^{(i)}\right\},$
and
$\mathcal{A}_{-}:=\left\{\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[\tilde{B}]_{lj}^{(i)}\leq\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}\leq 0\right\}.$
Using Fréchet inequalities stated in Theorem C.2, we get:
$\mathbb{P}\left[\mathcal{A}\right]\geq\max(\mathbb{P}\left[\mathcal{A}_{+}\right],\mathbb{P}\left[\mathcal{A}_{-}\right]).$
Under Assumptions 3.2 and 3.3, Lemma 3.1 gives:
$\hat{\gamma}_{(k,i)}\sim\mathcal{N}\left(\gamma_{(k,i)}^{*},\dfrac{\sigma^{2}}{pqT}\right),$ (15)
where:
$\hat{\gamma}_{(k,i)}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}\quad\text{and}\quad\gamma_{(k,i)}^{*}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[{\mathbb{C}}\right]_{(k,l),(i,j)}.$
Under the same assumptions, as detailed in the proof of Theorem 3.4:
$\hat{\beta}_{i}\sim\mathcal{N}\left(\beta^{*},\dfrac{\sigma^{2}m}{(n-1)pqT}\right),$ (16)
where:
$\hat{\beta}_{i}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\tilde{B}\right]_{lj}^{(i)}\quad\text{and}\quad\beta^{*}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[B^{*}\right]_{lj}.$
By construction, $\hat{\gamma}_{(k,i)}$ and $\hat{\beta}_{i}$ are independent. In addition, as $\gamma_{(k,i)}^{*}=[A^{*}]_{ik}\beta^{*}$, Assumption 3.5 ensures that their expected values satisfy:
$0<\gamma_{(k,i)}^{*}<\beta^{*}.$
Using the symmetry of the Gaussian distribution, we deduce:
$\mathbb{P}\left[\mathcal{A}_{+}\right]\geq\mathbb{P}\left[\mathcal{A}_{-}\right]\quad\text{and}\quad\mathbb{P}\left[\mathcal{A}\right]\geq\mathbb{P}\left[\mathcal{A}_{+}\right].$
The event $\mathcal{A}_{+}$ holds if $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ hold simultaneously, where:
$\mathcal{B}_{1}:=\left\{0\leq\hat{\gamma}_{(k,i)}\right\}\quad\text{and}\quad\mathcal{B}_{2}:=\left\{\hat{\gamma}_{(k,i)}\leq\hat{\beta}_{i}\right\}.$
Using Fréchet inequalities:
$\mathbb{P}\left[\mathcal{A}_{+}\right]\geq\mathbb{P}\left[\mathcal{B}_{1}\right]+\mathbb{P}\left[\mathcal{B}_{2}\right]-1.$
We now bound from above the probability of the events $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ separately. For $\mathcal{B}_{1}$, we note:
$\mathcal{B}_{1}=\left\{\gamma_{(k,i)}^{*}\geq\gamma_{(k,i)}^{*}-\hat{\gamma}_{(k,i)}\right\}.$
The complementary event $\bar{\mathcal{B}_{1}}$ is:
$\bar{\mathcal{B}_{1}}=\left\{\gamma_{(k,i)}^{*}\leq\gamma_{(k,i)}^{*}-\hat{\gamma}_{(k,i)}\right\}.$
As we have $\gamma_{(k,i)}^{*}>0$, using (15) and Mill’s inequality stated in Theorem C.1 provides:
$\mathbb{P}\left[\bar{\mathcal{B}_{1}}\right]\leq\frac{\sigma}{\gamma_{(k,i)}^{*}\sqrt{2pqT\pi}}\cdot\exp\left(-\dfrac{Tpq\left(\gamma_{(k,i)}^{*}\right)^{2}}{2\sigma^{2}}\right).$
For $\mathcal{B}_{2}$ we note that:
$\mathcal{B}_{2}=\left\{\beta^{*}-\gamma_{(k,i)}^{*}\geq(\beta^{*}-\gamma_{(k,i)}^{*})-(\hat{\beta}_{i}-\hat{\gamma}_{(k,i)})\right\}.$
The complementary event $\bar{\mathcal{B}_{2}}$ is:
$\bar{\mathcal{B}_{2}}=\left\{\beta^{*}-\gamma_{(k,i)}^{*}\leq(\beta^{*}-\gamma_{(k,i)}^{*})-(\hat{\beta}_{i}-\hat{\gamma}_{(k,i)})\right\}.$
As we have $\beta^{*}-\gamma_{(k,i)}^{*}>0$, using independence of $\hat{\gamma}_{(k,i)}$ and $\hat{\beta}_{i}$, results from (15), (16) and Mill’s inequality ensures:
$\mathbb{P}\left[\bar{\mathcal{B}_{2}}\right]\leq\frac{\sigma\sqrt{1+\frac{m}{n-1}}}{(\beta^{*}-\gamma_{(k,i)}^{*})\sqrt{2pqT\pi}}\cdot\exp\left(-\dfrac{Tpq\left(\beta^{*}-\gamma_{(k,i)}^{*}\right)^{2}}{2\sigma^{2}}\right).$
Finally, using:
$\mathbb{P}\left[\mathcal{A}_{+}\right]\geq 1-\mathbb{P}\left[\bar{\mathcal{B}_{1}}\right]-\mathbb{P}\left[\bar{\mathcal{B}_{2}}\right],$
we deduce the result. ∎
Proof of Theorem 3.7.
Let us fix $(i,k)\in[n]\times[m]$. We work on the event $\mathcal{A}:=\left\{[\tilde{A}]_{ik}=[\hat{A}]_{ik}\right\}$. Using (9), we get:
$\frac{1}{pq}[\tilde{A}]_{ik}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[\tilde{B}]_{lj}^{(i)}=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}.$
Proposition 2.4 guarantees:
$\frac{1}{pq}[A^{*}]_{ik}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}[B^{*}]_{lj}=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[{\mathbb{C}}\right]_{(k,l),(i,j)}.$
By definition of $\beta^{*}$ and using the reverse triangle’s inequality we get:
$\left\lvert\beta^{*}\right\rvert\cdot\left\lvert[\tilde{A}]_{ik}-[A^{*}]_{ik}\right\rvert\leq\left\lvert\hat{\gamma}_{(k,i)}-\gamma_{(k,i)}^{*}\right\rvert+\left\lvert[\tilde{A}]_{ik}\right\rvert\left\lvert\hat{\beta}_{i}-\beta^{*}\right\rvert,$
where:
$\hat{\beta}_{i}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\tilde{B}\right]_{lj}^{(i)},\;\hat{\gamma}_{(k,i)}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[\widehat{\mathbb{C}}\right]_{(k,l),(i,j)}$
and
$\gamma_{(k,i)}^{*}:=\frac{1}{pq}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}\left[{\mathbb{C}}\right]_{(k,l),(i,j)}.$
Under Assumptions 3.2 and 3.3, (15) and (16) hold. Using (15) and Mill’s inequality stated in Theorem C.1, we get for any $\epsilon>0$:
$\mathbb{P}\left[\left\lvert\hat{\gamma}_{(k,i)}-\gamma_{(k,i)}^{*}\right\rvert>\epsilon\right]\leq\frac{\sigma\sqrt{2}}{\epsilon\sqrt{pqT\pi}}\exp\left(-\frac{\epsilon^{2}pqT}{2\sigma^{2}}\right).$
Using (16) and Mill’s inequality, we find for any $\epsilon>0$:
$\mathbb{P}\left[\left\lvert\hat{\beta}_{i}-\beta^{*}\right\rvert>\epsilon\right]\leq\frac{\sigma\sqrt{2m}}{\epsilon\sqrt{pqT(n-1)\pi}}\exp\left(-\frac{\epsilon^{2}pqT(n-1)}{2m\sigma^{2}}\right).$
Finally, on the event $\mathcal{A}$, we have $\left\lvert[\tilde{A}]_{ik}\right\rvert\leq 1$. We then use the result from Lemma 3.6, Fréchet inequality stated in Theorem C.2, and we conclude with a union bound. ∎
Proof of Theorem 3.8.
For all $(l,j)\in[q]\times[p]$, (14) ensures that:
$[\hat{B}]_{lj}=[B^{*}]_{lj}+\epsilon_{lj},$
where:
$\epsilon_{lj}\overset{i.i.d.}{\sim}\mathcal{N}\left(0,\frac{m\sigma^{2}}{nT}\right).$
From the results in Appendix B, we obtain that the event $\mathcal{A}$ holds with probability at least $1-\delta$ where $\mathcal{A}$ is the event
$\max\limits_{\begin{subarray}{c}l\in[q]\\ j\in[p]\end{subarray}}\left|\epsilon_{lj}\right|\leq\sigma\sqrt{\frac{2m}{nT}}\left(\sqrt{\log(2pq)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right).$
We recall the definition of the threshold:
$\tau:=\sigma\sqrt{\frac{2m}{nT}}\left(\sqrt{\log(2pq)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right).$
On the event $\mathcal{A}$, we observe the following:
  • If $\lvert[\hat{B}]_{lj}\rvert>2\tau$, then:
    $\lvert[B^{*}]_{lj}\rvert\geq\lvert[\hat{B}]_{lj}\rvert-\lvert\epsilon_{lj}\rvert>\tau.$
  • If $\lvert[\hat{B}]_{lj}\rvert\leq 2\tau$, then:
    $\lvert[B^{*}]_{lj}\rvert\leq\lvert[\hat{B}]_{lj}\rvert+\lvert\epsilon_{lj}\rvert\leq 3\tau.$
From these observations, we deduce:
$\left\lvert[\hat{B}^{S}]_{lj}-[B^{*}]_{lj}\right\rvert=\begin{cases}\left\lvert[\hat{B}]_{lj}-[B^{*}]_{lj}\right\rvert,&\text{if }\lvert[\hat{B}]_{lj}\rvert>2\tau,\\ \lvert[B^{*}]_{lj}\rvert,&\text{if }\lvert[\hat{B}]_{lj}\rvert\leq 2\tau.\end{cases}$
Rewriting this equality with indicator functions, we have:
$\left\lvert[\hat{B}^{S}]_{lj}-[B^{*}]_{lj}\right\rvert=\lvert\epsilon_{lj}\rvert\mathbbm{1}\left(\lvert[\hat{B}]_{lj}\rvert>2\tau\right)+\lvert[B^{*}]_{lj}\rvert\mathbbm{1}\left(\lvert[\hat{B}]_{lj}\rvert\leq 2\tau\right).$
This leads to the bound:
$\left\lvert[\hat{B}^{S}]_{lj}-[B^{*}]_{lj}\right\rvert\leq\tau\mathbbm{1}\left(\lvert[B^{*}]_{lj}\rvert>\tau\right)+3\tau\mathbbm{1}\left(\lvert[B^{*}]_{lj}\rvert\leq 3\tau\right).$
Finally, we conclude:
$\left\lvert[\hat{B}^{S}]_{lj}-[B^{*}]_{lj}\right\rvert\leq 4\min\left(\tau,\lvert[B^{*}]_{lj}\rvert\right).$
By definition of the Frobenius norm, we obtain:
$\left\|\hat{B}^{S}-B^{*}\right\|_{F}^{2}\leq 16\left\|B^{*}\right\|_{0}\tau^{2}.$
This proves the first point.
For the second point, on the event $\mathcal{A}$ and under the stated assumption, we observe the following:
  • If $\lvert[B^{*}]_{lj}\rvert\neq 0$, then $\lvert[B^{*}]_{lj}\rvert>3\tau$. This provides the bound:
    $\lvert[\hat{B}]_{lj}\rvert=\lvert[{B}^{*}]_{lj}+\epsilon_{lj}\rvert>2\tau.$
    Hence $\lvert[\hat{B}^{S}]_{lj}\rvert\neq 0$.
  • If $\lvert[\hat{B}^{S}]_{lj}\rvert\neq 0$, then $\lvert[\hat{B}]_{lj}\rvert>2\tau$. This provides the bound:
    $\lvert[B^{*}]_{lj}\rvert\geq\lvert[\hat{B}]_{lj}\rvert-\lvert\epsilon_{lj}\rvert\geq\tau.$
    Hence $\lvert[{B}^{*}]_{lj}\rvert\neq 0$.
Proof of Theorem 3.9.
For all $(i,k)\in[n]\times[m]$, (15) ensures that:
$[\hat{\gamma}]_{ik}=[\gamma^{*}]_{ik}+\epsilon_{ik},$
where:
[…]
Proposition 2.4 ensures that
$\gamma_{ik}^{*}=[A^{*}]_{ik}\beta^{*}.$
From the results in Appendix B, we obtain that the event $\mathcal{A}$ holds with probability at least $1-\delta$ where $\mathcal{A}$ is the event
$\max\limits_{\begin{subarray}{c}i\in[n]\\ k\in[m]\end{subarray}}\left|\epsilon_{ik}\right|\leq\sigma\sqrt{\frac{2}{Tpq}}\left(\sqrt{\log(2nm)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right).$
We recall the definition of the threshold:
$\tau:=\sigma\sqrt{\frac{2}{Tpq}}\left(\sqrt{\log(2nm)}+\sqrt{\log\left(\frac{2}{\delta}\right)}\right).$
On the event $\mathcal{A}$, we observe the following:
  • If $\lvert\hat{\gamma}_{ik}\rvert>2\tau$, then:
    $\lvert\gamma^{*}_{ik}\rvert\geq\lvert\hat{\gamma}_{ik}\rvert-\lvert\epsilon_{ik}\rvert>\tau.$
    Thus we have $[A^{*}]_{ik}\neq 0$ and because $[A^{*}]_{ik}$ is bounded from above by $1$ we also have $\beta^{*}\geq\tau.$
  • If $\lvert\hat{\gamma}_{ik}\rvert\leq 2\tau$, then:
    $\lvert\gamma^{*}_{ik}\rvert\leq\lvert\hat{\gamma}_{ik}\rvert+\lvert\epsilon_{ik}\rvert\leq 3\tau.$
From these observations, we deduce:
$\left\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\right\rvert=\begin{cases}\left\lvert[\hat{A}]_{ik}-[A^{*}]_{ik}\right\rvert,&\text{if }\lvert\hat{\gamma}_{ik}\rvert>2\tau,\\ \lvert[A^{*}]_{ik}\rvert,&\text{if }\lvert\hat{\gamma}_{ik}\rvert\leq 2\tau.\end{cases}$
Rewriting this with indicator functions and using the previously stated implications, we have:
$\left\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\right\rvert\leq\left\lvert[\hat{A}]_{ik}-[A^{*}]_{ik}\right\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert>\tau\right)+\left\lvert[A^{*}]_{ik}\right\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert\leq 3\tau\right).$
Multiplying all terms by $\lvert\beta^{*}\rvert$ leads to the bound:
$\lvert\beta^{*}\rvert\cdot\lvert[\hat{A}^{S}]_{lj}-[A^{*}]_{lj}\rvert\leq\lvert\beta^{*}\rvert\left\lvert[\hat{A}]_{ik}-[A^{*}]_{ik}\right\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert>\tau\right)+\lvert\gamma^{*}_{ik}\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert\leq 3\tau\right).$
Using the equality
$\lvert\beta^{*}\rvert\left\lvert[\hat{A}]_{ik}-[A^{*}]_{ik}\right\rvert=\left\lvert(\beta^{*}-\hat{\beta})[\hat{A}]_{ik}+\hat{\beta}[\hat{A}]_{ik}-\beta^{*}[{A}^{*}]_{ik}\right\rvert$
and the triangle inequality leads to
$\lvert\beta^{*}\rvert\cdot\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\rvert\leq(M^{(1)}_{ik}+M^{(2)}_{ik})\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert>\tau\right)+\lvert\gamma^{*}_{ik}\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert\leq 3\tau\right),$
where
$M^{(1)}_{ik}=\lvert\beta^{*}-\hat{\beta}\rvert\left\lvert[\hat{A}]_{ik}\right\rvert\quad\text{and}\quad M^{(2)}_{ik}=\left\lvert\hat{\beta}[\hat{A}]_{ik}-\beta^{*}[{A}^{*}]_{ik}\right\rvert.$
Moreover we notice that
$M^{(2)}_{ik}=\left\lvert\hat{\gamma}_{ik}-\gamma_{ik}^{*}\right\rvert.$
Equations (15) and (16), together with Mill’s inequality, stated in Theorem C.1, provide for any $t>0$:
$\mathbb{P}\left[\left\lvert\beta^{*}-\hat{\beta}\right\rvert>t\right]\leq\frac{\sigma\sqrt{2m}}{t\sqrt{npqT\pi}}\exp\left(-\frac{Tpqt^{2}}{2\sigma^{2}m}\right)$
and
$\mathbb{P}\left[\left\lvert\gamma_{ik}^{*}-\hat{\gamma}_{ik}\right\rvert>t\right]\leq\frac{\sigma\sqrt{2}}{t\sqrt{pqT\pi}}\exp\left(-\frac{Tpqt^{2}}{2\sigma^{2}}\right).$
Finally, we get for any $\delta\in(0,1)$ with probability at least $1-\delta$:
$\lvert\beta^{*}\rvert\cdot\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\rvert\leq 2t_{\delta}\mathbbm{1}\left(A^{*}_{ik}\neq 0\right)\mathbbm{1}\left(\lvert\beta^{*}\rvert\geq\tau\right)+\lvert\gamma^{*}_{ik}\rvert\mathbbm{1}\left(\lvert\gamma^{*}_{ik}\rvert\leq 3\tau\right).$
Dividing by $\lvert\beta^{*}\rvert^{-1}$ and using that $\gamma^{*}_{ik}=[A^{*}]_{ik}\beta^{*}$ ensures that with probability at least $1-\delta$:
$\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\rvert\leq 2t_{\delta}\lvert\beta^{*}\rvert^{-1}\mathbbm{1}\left(A^{*}_{ik}\neq 0\right)\mathbbm{1}\left(\lvert\beta^{*}\rvert^{-1}\leq\tau^{-1}\right)+[A^{*}]_{ik}\mathbbm{1}\left([A^{*}]_{ik}\leq 3\tau\lvert\beta^{*}\rvert^{-1}\right).$
Finally we get with probability at least $1-\delta$:
$\lvert[\hat{A}^{S}]_{ik}-[A^{*}]_{ik}\rvert\leq$ $\lvert\beta^{*}\rvert^{-1}\mathbbm{1}\left(A^{*}_{ik}\neq 0\right)\left(2t_{\delta}+3\tau\right).$
By definition of the Frobenius norm, we obtain with probability at least $1-\delta$:
$\left\|\hat{A}^{S}-A^{*}\right\|_{F}^{2}\leq\lvert\beta^{*}\rvert^{-2}\left\|A^{*}\right\|_{0}\left(2t_{\delta}+3\tau\right)^{2}.$
For the second point, on the event $\mathcal{A}$ and under the stated assumptions, we observe the following:
  • If $\lvert[A^{*}]_{ik}\rvert\neq 0$, then $\lvert[A^{*}]_{ik}\rvert>3\tau$. Using the equality ${\gamma}^{*}_{ik}=\beta^{*}[A^{*}]_{ik}$ and the assumption on the bound satisfied by $\beta^{*}$ provide:
    $\lvert\hat{\gamma}_{ik}\rvert=\lvert{\gamma}^{*}_{ik}+\epsilon_{ik}\rvert>3\beta^{*}\tau-\tau>0.$
    Hence $\lvert[\hat{A}^{S}]_{ik}\rvert\neq 0$.
  • If $\lvert[\hat{A}^{S}]_{ik}\rvert\neq 0$, then $\lvert\hat{\gamma}_{ik}\rvert>2\tau$. This provides the bound:
    $\lvert\gamma^{*}_{lj}\rvert\geq\lvert\hat{\gamma}_{lj}\rvert-\lvert\epsilon_{lj}\rvert\geq\tau.$
    Hence $\lvert{\gamma}^{*}_{lj}\rvert\neq 0$ and thus $[A^{*}]_{ik}\neq 0$.
Appendix B Tails of the Maximum of Absolute Gaussian Variables
In this appendix, adapted from (Tu,, 2017), we analyze the upper and lower tails of the random variable $\max_{1\leq i\leq n}|X_{i}|$, where $X_{1},\ldots,X_{n}$ are independent and identically distributed (i.i.d.) random variables following a standard normal distribution $\mathcal{N}(0,1)$. Specifically, we aim to establish bounds for the upper and lower tails of this random variable with high probability, which is crucial for understanding the behavior of maxima in Gaussian settings.
We begin by stating the main result.
Theorem B.1.
Fix $\delta\in(0,1)$, and let $X_{1},\ldots,X_{n}$ be i.i.d. $\mathcal{N}(0,1)$. With probability at least $1-\delta$,
$\sqrt{\frac{\pi}{2}}\sqrt{\log(n/2)-\log\log(2/\delta)}\leq\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\sqrt{2}\left(\sqrt{\log(2n)}+\sqrt{\log(2/\delta)}\right).$
The asymmetry in the tails arises from the fact that $\max_{1\leq i\leq n}|X_{i}|$ is bounded below by zero almost surely, it is not bounded above by any fixed constant. To establish the theorem, we separately analyze the upper and lower tails and combine these results with a union bound.
B.1 Upper Tail
The upper tail is analyzed using concentration results for the suprema of Gaussian processes. The key tool is Talagrand’s concentration inequality, Lemma 2.10.6 in (Talagrand,, 2014), stated in a simpler version as follows:
Lemma B.2.
Consider $T\subseteq\mathbb{R}^{n}$ and let $g\sim\mathcal{N}(0,I)$. Then, for any $u>0$,
$\mathbb{P}\left\{\sup_{t\in T}\langle t,g\rangle-\mathbb{E}\left[\sup_{t\in T}\langle t,g\rangle\right]>u\right\}\leq\exp\left\{-\frac{u^{2}}{2s^{2}}\right\},$
where $s:=\sup_{t\in T}\mathbb{E}[\langle t,g\rangle^{2}]^{1/2}$.
To apply this inequality, we consider $T=\{e_{1},\ldots,e_{n},-e_{1},\ldots,-e_{n}\}$, where $e_{i}$ is the $i$-th canonical basis vector in $\mathbb{R}^{n}$. For $g\sim\mathcal{N}(0,I)$, this gives
$\sup_{t\in T}\langle t,g\rangle=\max_{1\leq i\leq n}\lvert g_{i}\rvert\quad\text{and}\quad\mathbb{E}[\langle t,g\rangle^{2}]=\mathbb{E}[g_{i}^{2}]=1.$
Applying Lemma B.2, we obtain, for any $\delta>0$:
$\mathbb{P}\left(\max_{1\leq i\leq n}\lvert X_{i}\rvert>\mathbb{E}\left[\max_{1\leq i\leq n}\lvert X_{i}\rvert\right]+\sqrt{2\log(1/\delta)}\right)\leq\delta.$
Thus, with probability at least $1-\delta$:
$\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\mathbb{E}\left[\max_{1\leq i\leq n}\lvert X_{i}\rvert\right]+\sqrt{2\log(1/\delta)}.$ (17)
The next step is to bound $\mathbb{E}\left[\max_{1\leq i\leq n}\lvert X_{i}\rvert\right]$ from above.
B.2 Expected Maximum of Gaussian Variables
We now bound $\mathbb{E}\left[\max_{1\leq i\leq n}\lvert X_{i}\rvert\right]$.
Proposition B.3.
Let $Z_{1},\ldots,Z_{n}$ be $n$ random variables (not necessarily independent) with marginal distribution $\mathcal{N}(0,1)$. Then,
$\mathbb{E}\left[\max_{1\leq i\leq n}Z_{i}\right]\leq\sqrt{2\log{n}}.$
Proof.
Fix any $\lambda>0$. Observe that
$\mathbb{E}\left[e^{\lambda\max_{1\leq i\leq n}Z_{i}}\right]\leq\sum_{i=1}^{n}\mathbb{E}\left[e^{\lambda Z_{i}}\right]=ne^{\lambda^{2}/2}.$
Taking the logarithm of both sides,
$\log\mathbb{E}\left[e^{\lambda\max_{1\leq i\leq n}Z_{i}}\right]\leq\log{n}+\frac{\lambda^{2}}{2}.$
Applying Jensen’s inequality,
$\lambda\mathbb{E}\left[\max_{1\leq i\leq n}Z_{i}\right]\leq\log\mathbb{E}\left[e^{\lambda\max_{1\leq i\leq n}Z_{i}}\right]\leq\log{n}+\frac{\lambda^{2}}{2}.$
Dividing through by $\lambda$,
$\mathbb{E}\left[\max_{1\leq i\leq n}Z_{i}\right]\leq\frac{\log{n}}{\lambda}+\frac{\lambda}{2}.$
Optimizing by setting $\lambda=\sqrt{2\log{n}}$, we conclude
$\mathbb{E}\left[\max_{1\leq i\leq n}Z_{i}\right]\leq\sqrt{2\log{n}}.$
For $X_{1},\ldots,X_{n}$ i.i.d. following a standard normal distribution $\mathcal{N}(0,1)$, considering the $2n$ variables $(X_{1},\ldots,X_{n},-X_{1},\ldots,-X_{n})$, we have:
$\mathbb{E}\left[\max_{1\leq i\leq n}\lvert X_{i}\rvert\right]\leq\sqrt{2\log(2n)}.$
Substituting this into (17) completes the proof of the upper tail.
B.3 Lower Bound
We now analyze the lower tail of $\max_{1\leq i\leq n}\lvert X_{i}\rvert$. Fix a positive $\tau>0$. Then,
$\Pr\left\{\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\tau\right\}=\Pr\left(\lvert X_{1}\rvert\leq\tau,\ldots,\lvert X_{n}\rvert\leq\tau\right).$
Using the independence of $X_{1},\ldots,X_{n}$, we get:
$\Pr\left\{\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\tau\right\}=\prod_{i=1}^{n}\Pr\left(\lvert X_{i}\rvert\leq\tau\right)$
The Gauss error function, defined as
[…]
allows then to write:
$\Pr\left\{\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\tau\right\}=\mathrm{erf}\left(\frac{\tau}{\sqrt{2}}\right)^{n}$
The inequality $\mathrm{erf}(x)^{2}\leq 1-e^{-4x^{2}/\pi}$, which holds for all $x\geq 0$, provides:
$\Pr\left\{\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\tau\right\}\leq\left(1-e^{-\frac{2}{\pi}\tau^{2}}\right)^{n/2}$
Finally, the inequality $1-x\leq e^{-x}$, which holds for all $x\in\mathbb{R}$ ensures:
$\Pr\left\{\max_{1\leq i\leq n}\lvert X_{i}\rvert\leq\tau\right\}\leq\exp\left(-\frac{n}{2}e^{-\frac{2}{\pi}\tau^{2}}\right).$
To derive the lower bound, set
$\exp\left(-\frac{n}{2}e^{-\frac{2}{\pi}\tau^{2}}\right)=\delta,$
and solve for $\tau$. We find that with probability at least $1-\delta$,
$\max_{1\leq i\leq n}\lvert X_{i}\rvert\geq\sqrt{\frac{\pi}{2}\log(n/2)-\frac{\pi}{2}\log\log(1/\delta)}.$
Appendix C Probability bounds and inequalities
C.1 Mill’s inequality
Theorem C.1 (Mill’s Inequality).
Let $X$ be a Gaussian random variable with mean $\mu$ and variance $\sigma^{2}$. Then, for any $t>0$, the following inequality holds:
$\mathbb{P}(X-\mu>t)\leq\frac{\sigma}{t\sqrt{2\pi}}e^{-\frac{t^{2}}{2\sigma^{2}}}.$
By symmetry, we also have:
$\mathbb{P}(X-\mu<-t)\leq\frac{\sigma}{t\sqrt{2\pi}}e^{-\frac{t^{2}}{2\sigma^{2}}}.$
Proof.
A proof of this theorem can be found in (Rigollet & Hütter,, 2023). Let $Z=\frac{X-\mu}{\sigma}$, so that $Z\sim\mathcal{N}(0,1)$. The survival function of $Z$ is given by:
$\mathbb{P}(X-\mu>t)=\mathbb{P}(Z>z),\quad\text{where }z=\frac{t}{\sigma}.$
The standard normal survival function is expressed as:
[…]
where $\varphi(u)=\frac{1}{\sqrt{2\pi}}e^{-u^{2}/2}$ is the probability density function of $Z$. We aim to bound this integral.
Since $\varphi(u)$ decreases exponentially as $u$ increases, the integral is dominated by the exponential term. For $u\geq z>0$, we have:
[…]
Substituting $z=\frac{t}{\sigma}$, the inequality becomes:
$\mathbb{P}(X-\mu>t)\leq\frac{\sigma}{t\sqrt{2\pi}}e^{-\frac{t^{2}}{2\sigma^{2}}}.$
By the symmetry of the normal distribution, the probability of the lower tail can be similarly bounded:
$\mathbb{P}(X-\mu<-t)=\mathbb{P}(Z<-z)=\mathbb{P}(Z>z),$
which yields:
$\mathbb{P}(X-\mu<-t)\leq\frac{\sigma}{t\sqrt{2\pi}}e^{-\frac{t^{2}}{2\sigma^{2}}}.$
C.2 Fréchet inequalities
Theorem C.2 (Fréchet inequalities).
Let $A$ and $B$ be two events. The probability of their intersection satisfies:
$\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)\leq\mathbb{P}(A\cap B)\leq\min(\mathbb{P}(A),\mathbb{P}(B)).$
The probability of their union satisfies:
$\max(\mathbb{P}(A),\mathbb{P}(B))\leq\mathbb{P}(A\cup B)\leq\min(1,\mathbb{P}(A)+\mathbb{P}(B)).$
Proof.
We first aim to prove:
$\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1)\leq\mathbb{P}(A\cap B)\leq\min(\mathbb{P}(A),\mathbb{P}(B)).$
By the definition of probability, the probability of the intersection of two events $A$ and $B$ cannot exceed the probability of either event:
$\mathbb{P}(A\cap B)\leq\mathbb{P}(A),\quad\mathbb{P}(A\cap B)\leq\mathbb{P}(B).$
Taking the minimum of the two bounds, we obtain:
$\mathbb{P}(A\cap B)\leq\min(\mathbb{P}(A),\mathbb{P}(B)).$
The probability of the union of $A$ and $B$ is given by the inclusion-exclusion principle:
$\mathbb{P}(A\cup B)=\mathbb{P}(A)+\mathbb{P}(B)-\mathbb{P}(A\cap B).$
Since probabilities are non-negative, we must have:
$\mathbb{P}(A\cup B)\leq 1\implies\mathbb{P}(A)+\mathbb{P}(B)-\mathbb{P}(A\cap B)\leq 1.$
Rearranging, we get:
$\mathbb{P}(A\cap B)\geq\mathbb{P}(A)+\mathbb{P}(B)-1.$
Additionally, probabilities are non-negative, so:
$\mathbb{P}(A\cap B)\geq 0.$
Combining these two conditions, we conclude:
$\mathbb{P}(A\cap B)\geq\max(0,\mathbb{P}(A)+\mathbb{P}(B)-1).$
This concludes.
We now prove:
$\max(\mathbb{P}(A),\mathbb{P}(B))\leq\mathbb{P}(A\cup B)\leq\min(1,\mathbb{P}(A)+\mathbb{P}(B)).$
The probability of the union of $A$ and $B$ must be at least as large as the probability of either event, since $A\cup B$ includes all outcomes in $A$ or $B$:
$\mathbb{P}(A\cup B)\geq\mathbb{P}(A),\quad\mathbb{P}(A\cup B)\geq\mathbb{P}(B).$
Taking the maximum of these bounds, we obtain:
$\mathbb{P}(A\cup B)\geq\max(\mathbb{P}(A),\mathbb{P}(B)).$
From the inclusion-exclusion principle, we have:
$\mathbb{P}(A\cup B)=\mathbb{P}(A)+\mathbb{P}(B)-\mathbb{P}(A\cap B).$
Since $\mathbb{P}(A\cap B)\geq 0$, it follows that:
$\mathbb{P}(A\cup B)\leq\mathbb{P}(A)+\mathbb{P}(B).$
Additionally, probabilities cannot exceed 1, so:
$\mathbb{P}(A\cup B)\leq 1.$
Combining these two conditions, we conclude:
$\mathbb{P}(A\cup B)\leq\min(1,\mathbb{P}(A)+\mathbb{P}(B)).$
Appendix D Additional numerical simulations
In this section, we extend the numerical simulations presented in the main text by evaluating the performance of the proposed estimators $\hat{A}$ and $\hat{B}$ under the operator norm and max norm metrics. These additional results complement the Frobenius norm analysis, providing a more comprehensive assessment of the estimators’ performance.
Figures 5 and 6 present the error variation of the estimators $\hat{A}$ and $\hat{B}$ under the operator norm as functions of the parameters $n$, $m$, $p$, $q$, and $T$.
(a) EV w.r.t $n$
(b) EV w.r.t $m$
(c) EV w.r.t $T$
(d) EV w.r.t $p$
(e) EV w.r.t $q$
Figure 5: Error variation (EV) of the operator norm errors of $\hat{A}$ for estimating $A^{*}$ with respect to (w.r.t.) different parameters
(a) EV w.r.t $n$
(b) EV w.r.t $m$
(c) EV w.r.t $T$
(d) EV w.r.t $p$
(e) EV w.r.t $q$
Figure 6: Error variation (EV) of the operator norm errors of $\hat{B}$ for estimating $B^{*}$ with respect to (w.r.t.) different parameters
Figures 7 and 8 present the error variation of $\hat{A}$ and $\hat{B}$ under the max norm.
(a) EV w.r.t $n$
(b) EV w.r.t $m$
(c) EV w.r.t $T$
(d) EV w.r.t $p$
(e) EV w.r.t $q$
Figure 7: Error variation (EV) of the maximum norm errors of $\hat{A}$ for estimating $A^{*}$ with respect to (w.r.t.) different parameters
(a) EV w.r.t $n$
(b) EV w.r.t $m$
(c) EV w.r.t $T$
(d) EV w.r.t $p$
(e) EV w.r.t $q$
Figure 8: Error variation (EV) of the maximum norm errors of $\hat{B}$ for estimating $B^{*}$ with respect to (w.r.t.) different parameters
From the results, we observe similar trends in error variation as functions of the parameters $n,m,p,q,T$ under the operator norm and max norm as observed for the Frobenius norm in the main text. Specifically, the accuracy of the estimators improves with an increase in sample size $T$, while the effects of other parameters such as $n,m,p,q$ are consistent with the theoretical findings detailed in Theorems 3.7 and 3.4.
These additional results reinforce the theoretical properties of the proposed estimators under various norm-based metrics, providing a comprehensive understanding of their behavior across different settings.